A Complete Bibliography of Publications in Numerical Algebra, Control and Optimization

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

06 April 2022 Version 1.04

Title word cross-reference

(F, E) [150]. (ρ, η, A) [24]. ∗ [135]. 1 [49]. 3 [14]. A + BXB^{∗} [205]. α [147]. AXA^{∗} = B [305]. AXB = C [382]. Γ [245]. S [245]. C_{k} [49]. C_{m} [49]. E [439, 152]. f_{i} [146]. G [150]. H^{1} [16]. H_{∞} [427, 182, 360]. i ≠ j [76]. l^{k} [28]. L_{1} [197]. L_{∞} [279]. M [143]. R^{n} [128]. P_{∗} (κ) [407]. ϕ [174]. S [157]. SE(3) [321]. θ [336]. V [439]. V_{i}V_{j}^{T} = O [76].

* [135]. *-representation [135].

-algebras [135]. -Benson [152]. -constrained [197]. -convergence [245]. -differentiable [439]. -eigenvalues [28]. -expansive [146]. -goodness [157]. -HLCP [407]. -homogenization [245]. -invexity [439, 24]. -Lyapunov [16]. -matrix [143]. -norm [279]. -uniformly [174]. -well-posedness [147].

2D [241, 182]. 2D-delayed [182]. 2nd [260]. 2nd-order [260].

a-posteriori [85]. absolute [340, 278, 421, 403]. accelerated [249]. access [48]. achievement [58]. Active [15, 395, 427]. actuated [387]. actuator [378]. actuators [82]. Adaptive [248, 259, 391, 132, 140]. Adding [364]. adequate [289]. Adjacent [159, 32]. adjoint [252]. adjusting [183]. adjustment [113, 214]. Admission [45]. adult [399]. advances [3]. advection [226]. advective [391]. after [46]. after-call [46]. against [378]. agent [324, 264]. AGWO [411]. AGWO-PSO [411]. aided [351]. AIDS [330]. AIS [58]. AIS-based [58]. algebraic [316, 237, 167, 143]. algebras [158, 135]. algorithm [233, 271, 314, 79, 369, 65, 371, 242, 411, 30, 301, 440, 317, 56, 11, 9, 14, 161, 383, 116, 202, 299, 149, 287, 54, 169, 192, 373, 188, 437, 430, 36, 138, 29, 10, 351, 270, 63, 124, 195]. algorithmic [39]. algorithms [199, 144, 134, 60, 184, 326, 190, 322]. Allen [391]. allocation [356, 347]. alternate [415]. Alternating [372, 126, 396]. Analysis [48, 184, 263, 79, 21, 6, 242, 330, 132, 390, 410, 22, 169, 300, 365, 333, 241, 351, 85, 191, 346]. Analytical [47]. Analyzing [98]. Angola [90]. ant [394]. Anti [276, 399]. Anti-gene [276]. anti-predator [399]. AOR [432, 103]. Application [317, 402, 137, 66, 213, 99, 261, 225, 94, 316]. applications [431, 135, 440, 142, 175, 428]. applied [291, 362, 171]. Applying [434]. Approach [339, 62, 338, 66, 110, 35, 148, 172, 409, 43, 381, 194, 224, 384, 302, 8]. approaches [154, 280, 49]. approximant [434]. Approximate [293, 404, 292, 175]. approximating [430]. Approximation [144, 88, 344, 4, 246, 406, 240, 190]. approximations [34]. arbitrary [54]. arc [287]. arc-search [287]. architectured [252]. arcwise [215]. arising [86]. arm [62, 160]. ARMA [329, 362]. artificial [65, 410, 413]. aspects [39, 440]. assignment [123]. associated [421]. Asymptotic [38, 285, 406, 258]. asymptotically [146]. Asynchronous [42]. augmented [35, 103, 216]. automata [206]. automatic [59]. autonomous [168, 404]. Auxiliary [167]. average [200, 419]. averaging [258]. averse [374]. avoidance [84]. Axiomatic [418].

back [126]. backward [130, 248].

Balancing [213]. ball [387, 350]. Banach [277, 39, 344, 174, 404, 18]. bandwidth [45]. bang [88, 87, 381]. bang-bang [88, 381]. bang-singular [87]. barrier [290, 412, 191]. based [434, 213, 58, 4, 318, 13, 353, 407, 9, 14, 218, 154, 249, 116, 145, 149, 410, 160, 226, 234, 43, 184, 373, 117, 194, 385, 306, 388, 347, 255, 324, 206, 163, 191, 322, 270, 360, 216, 339, 267, 186, 8, 201]. batch [207, 187, 210, 212]. Baxter [133]. beamformer [231]. Bearing [321]. bee [65]. Behavior [398, 418, 399]. Benard [332]. benchmark [395]. Benson [152]. bent [377]. best [165]. between [405, 258, 413]. bi [408]. bi-criterion [408]. bicycle [168]. bidirectional [45]. Bifurcation [365]. bike [420]. bilateral [357]. Bilevel [39, 62, 66, 134]. binary [332]. biological [163, 186]. biometrics [410]. bipedal [13]. bisection [195]. bisimulation [232]. bit [48]. block [103, 248]. blocking [54]. blood [266]. bodies [321]. boundary [283, 181, 349, 263, 57]. bounds [88, 74, 284, 219, 151, 308]. box [138, 368, 162]. box-constrained [368]. Bramble [105]. bridge [395]. Bridging [258]. brief [304]. broadcasting [42]. Broyden [37]. Broyden-like [37]. BSDEs [73]. bullwhip [153]. bundles [41]. Buongiorno [333].

Cahn [391]. calculus [119, 114]. call [46]. captured [319]. Caputo [438]. CAR [276]. Carathéodory [119]. cardinality [265]. Carlo [267]. Cartesian [159]. case [90, 302]. cash [431]. Celis [8]. cell [276]. cells [352]. cellular [206]. censored [249]. center [310]. certain [437]. Chain [73, 43, 281, 312, 389, 266]. changing [235]. Characterization [112, 345, 53, 50]. Characterizations [152]. cheap [285]. Chebyshev [294, 142, 438]. Chimeric [276]. Cholesky [286]. circular [199, 253]. class [61, 331, 338, 66, 294, 336, 407, 238, 188, 437, 364, 256, 433, 345, 150, 17, 64, 57]. climbing [30]. Closed [238, 389]. Closed-form [238]. closed-loop [389]. clusters [41]. code [138]. coding [42]. coefficient [214]. coefficients [80, 341]. coincide [357]. collinear [254]. collision [84]. colony [394, 65]. colorings [159]. combination [398, 322]. common [356, 146, 430]. communication [43, 360]. Comparison [405]. comparisons [340]. complementarity [396, 5, 166, 192, 300, 151, 367, 64]. Complete [128, 273, 267]. completion [6]. Complex [355, 261, 426, 325, 337, 106]. Complexity [191, 184, 231]. component [381]. component-wise [381]. components [98]. composite [156]. computation [279]. Computational [173, 252, 98, 342, 289, 235]. computer [351]. Computing [250, 107, 179, 371]. concept [232]. concurrent [232]. condition [77, 162, 310]. conditional [389]. Conditions [393, 61, 87, 118, 424, 26, 436, 256, 255, 24]. conductivity [185]. cone [199, 313, 166, 165, 37, 151, 421, 215, 55]. cone-arcwise [215]. Conformal [331]. conic [124, 131]. conjugate [315, 297]. connected [215]. connection [52]. consensus [327, 322, 264]. consensus-based [322]. consideration [164]. considering [312, 389]. consistent [384]. Constrained [327, 394, 265, 100, 25, 147, 94, 427, 409, 37, 412, 224, 256, 345, 139, 197, 368]. constraint [283]. constraints [298, 431, 56, 5, 313, 427, 292, 192, 240, 300, 89, 138, 350, 97, 415, 272, 55, 358, 8]. Construction [289, 308]. consuming [275]. contact [61]. container [154]. Continuity [198, 32, 175]. continuous [198, 15, 89, 436, 209, 194, 413, 419, 186]. Control [13, 89, 82, 414, 338, 352, 88, 283, 431, 79, 144, 19, 213, 342, 392, 239, 58, 397, 91, 207, 357, 395, 285, 181, 406, 427, 429, 69, 116, 202, 204, 311, 348, 245, 381, 361, 386, 168, 222, 438, 117, 259, 140, 385, 45, 119, 364, 92, 210, 90, 256, 111, 324, 137, 345, 272, 251, 354, 85, 86, 182, 235, 360, 153, 263, 17, 378]. Controllability [435, 261, 393, 409, 404]. controllable [187]. controlled [291, 10]. controller [433]. Controlling [47]. controls [357, 255]. Convergence [432, 6, 242, 370, 169, 107, 134, 108, 56, 5, 422, 245, 27, 146, 165, 174, 89, 183, 384, 216, 310]. convex [277, 199, 25, 242, 74, 20, 372, 126, 146, 188, 155, 322, 150, 376, 359, 63, 156]. convexity [228]. coordination [324]. correcting [196]. correction [340]. corrector [314]. correlation [102]. cost [62, 200, 244, 282, 190, 235]. costs [113]. countably [146]. Cournot [374]. cover [190]. coverage [111]. credit [282, 343]. criterion [408]. cross [266]. cross-efficiency [266]. crowdsourced [420]. CTL [352]. cubic [138, 379]. culture [207, 210, 186]. cure [330]. curvature [36]. curved [290]. curvilinear [226, 245]. CVaR [4]. CVaR-based [4]. cyclic [299]. cytokine [276].

D [14]. D-structure [14]. DAE [239]. damped [112]. Dampening [153]. damping [117]. data [423, 249, 424]. DC [66, 156]. DEA [267]. decentralised [232]. Decentralized [233]. decomposition [193, 216]. Decoupling [379]. definite [104, 6, 305, 154, 63]. Deflating [143]. Deflation [217]. deformations [331]. Delay [182, 91, 393, 41, 348, 363, 47, 438, 123, 51, 354]. Delay-range [182]. delayed [163, 182]. delays [392, 361, 365, 140]. delivery [41, 80]. demand [244, 343, 44]. Dennis [8]. density [384]. dependent [332, 277, 399, 244, 210, 182]. depots [268]. derivative [7, 438, 10, 375, 196]. derivative-free [7, 10, 375, 196]. Derivatives [221, 32]. descent [292, 293]. descriptor [213, 279, 218]. Design [394, 281, 231, 389, 82, 168, 222, 194, 167]. detection [15, 230, 167]. deteriorating [244, 343, 425]. Determining [181]. Development [232, 269]. deviation [278]. diagonally [284]. difference [177]. differences [69, 116, 202, 204]. differentiable [439, 313]. Differential [295, 434, 115, 400, 243, 225, 406, 316, 247, 252, 260, 59, 413, 334, 167, 349]. differential-algebraic [167]. differential-algebraic-equations [316]. differentiation [248, 47]. diffusion [336, 236, 226, 80, 177, 263]. diffusion-advection-reaction [226]. diffusion-type [336]. diffusive [332]. digital [325]. dilution [186]. dimension [264]. dimensional [31, 6, 223, 336, 226, 278, 80, 263]. Direct [414, 140, 86]. direction [372, 126, 396]. directions [315, 36]. discontinuous [391]. Discrete [433, 69, 116, 204, 234, 232, 436]. Discrete-time [433, 69, 116, 204, 234]. Discriminant [390]. disordering [362]. Dissipative [386]. distinguishing [159]. Distributed [326, 409, 324, 322]. distribution [94, 43]. Distributionally [428]. disturbances [285]. disturbed [363]. dominance [240, 203, 97]. dominant [284]. double [332, 128]. double-well [128]. Douglas [370]. down [290]. down-and-out [290]. Drazin [303]. driven [99, 53]. drug [80]. DTNs [41]. dual [199, 407, 184, 231, 36, 241, 191, 190]. duality [31, 38, 426, 424, 26, 170, 156]. due [113]. Duration [78]. Dynamic [206, 271, 418, 209, 45, 420, 193]. dynamical [327, 61, 431, 112, 163, 212]. Dynamics [399, 330].

echelon [281, 244]. edge [159]. edge-colorings [159]. effect [330, 333, 153]. Effective [246, 80]. efficiency [423, 266, 215, 152]. efficient [226, 345, 176, 267, 63]. eigensolvers [279]. eigenvalue [161, 37, 297, 123, 416, 367]. eigenvalues [371, 284, 148, 28, 221]. electric [172]. elements [29, 391]. emissions [194]. energy [172]. entropy [125, 266, 72]. environment [280, 211]. epidemic [206]. equality [292]. equation [434, 233, 283, 133, 405, 305, 382, 353, 406, 421, 404, 403, 349, 391]. equations [400, 432, 16, 246, 294, 336, 141, 132, 225, 316, 247, 250, 340, 145, 237, 7, 260, 413, 177, 334, 337, 167, 143, 106, 3]. equilibrium [39, 344, 134, 243, 422, 388, 55, 374]. Error [88, 151, 10, 85]. Essential [179]. estimates [94]. estimating [80]. estimation [369, 363]. estimations [303]. estimator [278]. Euclidean [158, 393, 162]. Euler [88, 16]. evaluation [52, 266, 12]. even [148]. event [232]. evidential [269]. evolution [295, 435, 59, 404]. exact [224]. Examination [392]. example [11]. exchanges [78]. excitation [13]. excited [395]. Existence [344, 147, 165, 39]. exits [119]. expansion [234]. expansive [146]. experiment [394]. Experiments [286]. exploiting [36]. exponentially [290, 309]. expression [238]. extension [287]. extensions [197]. Extragradient [25, 437]. Extragradient-projection [25]. extremal [255]. extremals [87]. extreme [371, 261]. extremely [226].

faces [230]. factorization [171].

factorizations [21]. factors [269]. failure [249, 167]. failures [378]. fair [48]. fairness [48]. family [71, 146, 174]. Fan [101]. Fault [363, 378, 15]. Fault-tolerant [378]. faults [15]. feasibility [242, 18]. feature [383]. fed [207, 210]. fed-batch [207, 210]. Feedback [256, 323, 16, 48, 222, 266, 92]. feeds [207]. Fenchel [156]. fermentation [209, 163, 211, 212]. field [384]. filling [11]. Filter [9, 192, 68]. Filter-based [9]. Filtering [69]. filters [325]. finance [99]. financial [312]. Finding [29, 148]. finite [174, 177, 29, 391]. fir [325]. first [405]. fish [365]. fitted [309]. fixed [146, 430]. fixing [419]. Fletcher [7]. flexible [230]. flow [271, 213, 179, 286, 312, 333, 351]. fluid [332, 44]. following [440, 166, 117]. force [160]. forcing [231]. form [225, 238, 252]. Formal [351]. formation [321]. forms [49]. formulas [353, 248, 255]. formulation [4]. forward [130]. forward-backward [130]. Forword [227]. four [266]. four-stage [266]. Fourier [274, 377]. Fourier-splitting [274]. fourth [309]. fraction [234]. fractional [260, 26, 438, 408, 307, 177, 349, 433, 24, 358, 8, 378, 201]. fractional-order [433]. frames [221]. framework [265, 58]. Fredholm [246]. free [7, 192, 217, 112, 10, 208, 375, 196]. Frequency [325, 234]. friction [16]. Frobenius [125]. frontal [230]. fuel [319]. fuel-minimal [319]. Fukushima [5]. Fukushima-Regularization [5]. full [149, 253, 216]. function [158, 16, 313, 383, 149, 184, 192, 205, 191, 377]. functional [399, 255]. functions [62, 6, 125, 75, 74, 200, 407, 94, 340, 155, 401, 224, 34, 150, 156]. Further [303]. Fused [278]. fusion [330]. Fuzzy [280, 259, 395, 361, 418, 362]. fuzzy-regression [280].

Galerkin [306, 391]. Galerkin-type [306]. game [110, 243, 326]. game-theoretic [110]. games [115]. gap [258, 308]. Gauss [376]. Gaussian [126, 94, 188, 211]. GE [435]. GE-evolution [435]. gene [276]. General [24, 225, 351, 139, 334]. generalization [104]. Generalized [388, 344, 369, 101, 103, 108, 257, 250, 422, 373, 26, 335, 24, 150, 55, 18]. generated [71]. generating [72]. generation [52, 326]. Genetic [14, 9, 29, 212]. geometric [60, 92, 12]. geometry [196]. given [162]. Global [257, 56, 409, 424, 59, 215, 223, 275, 60, 183]. Globalizer [275]. globally [96, 176]. glycerol [212]. goodness [157]. GPOPS [392]. GPOPS-II [392]. Gradient [94, 233, 202, 231, 297, 296, 324, 415, 81, 367]. gradient-based [324]. graphs [159]. Grasping [160]. greatest [292, 293]. green [281]. grey [262]. grid [394, 326]. grids [261]. group [203]. growing [273]. growth [113]. GSOR [380]. guess [342].

H [251]. H-Optimal [251]. Hadamard [74, 155]. Hahn [114]. Hamiltonian [394]. hand [160]. hand-arm [160]. hard [318]. Hardy [71]. hazardous [328]. Henig [176]. Hermite [74, 311, 348, 155]. Hermitian [104, 305, 382]. Hessian [324]. heuristic [271]. hidden [228]. high [223, 278]. higher [239, 401]. highway [395]. HIV [291, 352, 330]. HIV/AIDS [330]. HLCP [407]. Hölder [175]. holding [244, 282]. holonomy [331]. HOMO [308]. HOMO-LUMO [308]. Homogenization [245, 226, 258, 252]. Homotopy [318, 294]. horizon [118, 243, 285, 137]. housing [302]. HSS [310]. human [62]. hybrid [338, 271, 411, 181, 422, 386, 209, 413]. hyperbolic [274, 289]. hypersingular [294]. HZ [398].

ice [185]. idempotent [133]. Identification [65, 324, 209, 163, 212]. IER [266]. if [76]. II [392]. image [380, 410, 86, 170, 339]. immunology [291]. impact [98]. Imperfection [343]. implementation [286]. implicit [435, 174, 309]. important [269]. imprecise [423]. Improved [5, 295, 369, 30, 362]. improvement [347, 346]. Improving [383]. impulsive [19, 386, 210, 256]. incidence [171]. incipient [15]. inclusions [335]. increment [255]. Incremental [120, 336]. indefinite [372, 228]. Index [220, 136, 213, 239, 316, 48]. index-2 [213]. Index-proper [220, 136]. Index-range [136]. indexes [418]. Indirect [319].

Individual [410]. industry [402]. inequalities [158, 4, 147, 241, 216, 374]. inequality [101, 71, 75, 74, 89, 155, 437, 430, 224, 18]. inertia [205]. inertial [430]. inertial-projection [430]. inertias [382]. inexact [396, 416]. infeasible [340, 149, 253]. infected [352]. infection [397]. inference [249]. infinite [31, 118, 243, 285, 247, 146, 137]. infinite-dimensional [31]. Information [236, 440]. Initial [342]. initially [251]. inner [107]. input [393, 427, 361, 222, 140, 354]. input/output [140]. inspection [343]. inspired [256]. instabilities [332]. Instability [113]. Integer [97, 164, 262]. integral [283, 405, 246, 294, 433]. Integrated [312, 95, 282]. integration [260, 270]. integrator [364]. integro [334]. integro-differential [334]. intelligent [79]. intensity [50]. interconnected [222]. interior [314, 199, 407, 149, 253, 287, 184, 191]. interior-point [314, 199, 149, 287, 184, 191]. intermediate [431, 268]. internal [387]. Internet [52]. interpretation [92]. interval [432, 325, 249, 341, 339]. Introduction [142]. invariant [436]. invasive [317]. inventory [244, 282, 343, 153]. inverse [67, 238, 217, 171, 303]. inverse-free [217]. inversion [351]. investigation [269]. invexity [439, 24]. involving [207, 26]. irreducible [331, 143]. isolated [41]. isothermal [16]. issues [179]. items [244, 425]. iteration [107, 380, 104, 27, 174, 337, 384]. iterations [127]. Iterative [316, 279, 344, 108, 11, 116, 145, 373, 188, 349].

Jacobian [216, 106]. Jamming [110]. Jensen [75]. jobs [271, 187]. Jordan [158, 221]. jump [386]. jumps [47].

Kalman [436]. Karp [273]. kernel

[407, 149, 184, 191]. kind [405, 396, 373]. kinematics [168]. knapsack [419]. known [285]. Kronecker [49]. Krylov [217]. Kutta [309]. Ky [101].

Lagrange [156]. Lagrangian [35, 77, 216]. Lagrangian-based [216]. landing [247]. landmarks [230]. large [279, 319, 316, 237, 226, 416, 296, 81, 360]. large-scale [279, 237, 81, 360]. largest [161]. LASSO [278]. latently [352]. Laurent [219]. law [251]. layer [332, 214]. layout [45]. LCFS [50]. LCFS-PR [50]. leadership [115]. least [369, 305, 382, 373, 278, 304, 368, 3]. least-squares [304]. Legendre [438, 270]. levels [282]. Levenberg [421, 183]. Lévy [99]. libration [254]. lifetime [53]. like [336, 37, 416, 105]. Lin [5]. line [381]. Linear [115, 348, 334, 414, 88, 233, 314, 19, 31, 432, 103, 108, 30, 246, 243, 285, 229, 406, 370, 409, 250, 145, 166, 184, 297, 278, 82, 436, 112, 140, 307, 337, 256, 354, 85, 190, 68, 64, 142]. linear-quadratic [414, 88, 243, 406, 85]. linear/submodular [190]. Linearized [126]. Linearly [273]. Linearly-growing [273]. lines [51]. Lipschitz [60, 174, 310]. Littlewood [71]. Load [43]. Local [307, 310]. localization [328, 154]. localizing [137]. log [328]. Long [440]. Long-step [440]. longevity [58]. loop [115, 389]. Lorentzian [331]. low [231]. low-complexity [231]. lower [308]. LQ [228]. LQR [274]. LUMO [308]. Lyapunov [16, 316, 413].

machine [271, 187]. machines [12]. magnitudes [219]. makespan [187]. Malfatti [223, 257, 388]. management [431, 172]. manifolds [331]. manipulation [160]. MAPLE [138]. mappings [130, 74, 422, 146, 174, 430, 178]. maps [32, 215]. Marangoni [332]. margins [12]. market [302]. markets [302]. Markov [73, 43]. Markovian [73, 53, 386]. Marquardt [421, 183]. masses [387]. material [312]. Mathematical [135, 276, 5, 192, 300, 208]. matrices [76, 220, 238, 136, 171, 438, 106]. Matrix [203, 298, 133, 6, 135, 318, 305, 382, 49, 241, 324, 205, 351, 384, 143, 193, 379, 377]. matrix-valued [205]. max [195]. max-bisection [195]. maximal [71, 130, 102]. maximization [214]. maximizing [12]. Maximum [382, 72, 125, 119, 34]. ME [54]. Mead [412]. mean [41]. means [71]. measures [266]. Measuring [423]. mechanical [139]. mechanism [295]. mechanisms [54]. medium [336]. Mehrotra [314]. memetic [195]. mesh [142]. meta [271]. meta-heuristic [271]. metaheuristic [301]. method [394, 414, 107, 380, 400, 25, 344, 432, 158, 4, 103, 274, 130, 125, 405, 318, 246, 294, 395, 20, 372, 132, 225, 13, 407, 126, 396, 5, 370, 313, 250, 166, 253, 237, 292, 7, 37, 231, 297, 217, 102, 122, 416, 421, 385, 412, 413, 183, 337, 349, 415, 419, 81, 367, 272, 86, 208, 68, 153, 129, 216, 376, 55, 309, 57, 196, 131, 310, 201, 334]. methodology [408]. methods [319, 108, 398, 141, 14, 316, 145, 117, 304, 80, 72, 33, 403, 296, 2, 191, 170, 106, 3, 64]. metric [372, 27, 146, 165, 17]. MHSS [106]. micro [252]. micro-architectured [252]. microbial [207, 209, 163, 211, 186]. microgrids [172]. microstructure [226, 245]. middleware [79]. MIMO [231]. minimal [319, 29]. Minimax [178, 26, 262]. minimization [25, 318, 188]. minimize [271, 41, 187]. Minimum [229, 382, 250, 34]. mining [402]. MIQCP [98]. Missed [119]. mitigation [411]. Mixed [164, 115, 39, 344, 9, 54, 335, 18]. mobile [110]. mode [361]. Model [85, 291, 352, 356, 213, 164, 65, 397, 330, 218, 316, 325, 249, 69, 116, 202, 204, 234, 389, 365, 276, 333, 140, 399, 244, 282, 306, 343, 44, 206, 211, 302, 124, 201]. model-reality [69, 116, 202, 204]. Modeling [209, 212, 312, 171]. models [173, 99, 289, 267]. modification [130]. modified [313, 7, 437, 412, 296]. molecular [14]. monotone [130, 437, 34]. monotonicity [344, 136, 203]. Monte [267]. most [269]. motion [112, 254]. movements [62]. MRI [193]. Multi [296, 58, 426, 34, 324, 268, 264]. multi-agent [324, 264]. multi-objective [426, 34]. multi-robot [58]. Multi-step [296]. multi-trip [268]. multiclass [12]. multidimensional [390, 122]. multifingered [160]. multigrid [102]. multimedia [45]. multiobjective [439, 295, 138, 12]. multiple [321, 15, 207, 365, 42, 78, 24]. Multiplicative [21]. multipliers [372, 126, 396]. Multiserver [46]. multisplitting [108]. multistage [265]. multiuser [231]. multivalued [430].

nanofluid [333]. Nash [388, 55, 374]. navigator [45]. NDEA [266]. Necessary [118, 256]. negative [305, 36]. neighborhood [254]. Nelder [412]. Nesterov [149]. network [411, 53, 258, 281, 293, 312, 389, 171, 42, 194, 360]. networked [323]. networks [394, 327, 110, 261, 96, 229, 160, 226, 252, 245, 280, 43, 169, 45]. neural [160, 169, 293]. Neuro [395, 259]. Neuro-Fuzzy [259, 395]. neutrosophic [339]. Newton [6, 141, 353, 250, 416, 106, 55, 310, 201]. Newton-HSS [310]. Newton-like [416]. Newton-MHSS [106]. next [52]. node [53, 43]. Noether [91]. noise [211]. non [331, 145, 404, 359]. non-autonomous [404]. non-convex [359]. non-irreducible [331]. non-linear [145]. non-smooth [359]. nonconvergent [11]. nonconvex [35, 128, 129]. nonhomogeneous [161]. Nonlinear [175, 364, 210, 338, 400, 35, 77, 20, 225, 56, 396, 69, 116, 204, 292, 7, 304, 36, 413, 335, 2, 349, 137, 139, 163, 346, 68, 106, 3, 57]. Nonmonotone [131, 81, 367]. nonnegative [220, 161]. nonregular [77]. nonsmooth [188, 26]. nonstationary [43]. nonsymmetric [416]. norm [279, 250]. normal [328]. note [101, 145, 177, 34]. notion [203]. Novel [393, 411, 301, 275]. NP [273]. NP-complete [273]. NT [253]. NT-step [253]. number [271, 219, 122]. Numerical [330, 357, 340, 80, 387, 334, 272, 341, 440, 429, 260, 270, 3, 57]. Nutrient [365]. Nyström [309].

objective [426, 313, 34, 24]. observability [435]. observer [360]. observer-based [360]. obstacle [357, 341]. obstacles [357]. obtain [62]. ODEs [248]. on-demand [44]. one [48, 260, 193]. one-bit [48]. one-point [260]. Onset [332]. open [115, 54]. open-loop [115]. operation [96]. operational [438]. operator [435]. operators [125, 252]. Optimal [352, 323, 431, 397, 207, 41, 90, 251, 186, 214, 425, 327, 62, 414, 338, 283, 144, 19, 342, 392, 239, 58, 179, 91, 357, 96, 406, 142, 427, 69, 116, 202, 204, 424, 340, 311, 348, 245, 89, 438, 385, 119, 210, 254, 137, 272, 85, 86, 235, 153, 263, 17]. Optimality [426, 26, 185, 118, 77, 424, 51, 256, 255, 24]. Optimization [298, 394, 434, 265, 62, 100, 199, 31, 35, 398, 301, 223, 257, 275, 395, 317, 20, 372, 148, 56, 13, 94, 313, 383, 424, 149, 253, 287, 95, 60, 292, 381, 184, 293, 312, 363, 389, 128, 326, 59, 138, 33, 412, 413, 224, 350, 355, 296, 2, 415, 10, 81, 205, 32, 191, 375, 197, 359, 215, 358, 63, 124, 178, 162, 196, 152, 131, 428]. optimized [395]. optimizer [301, 262]. Optimum [328]. options [290]. Orbital [33, 254]. orbiters [319]. order [87, 148, 406, 247, 248, 313, 41, 260, 240, 436, 421, 401, 306, 177, 433, 32, 153, 55, 309, 57, 378]. order-up-to [153]. ordered [335, 29]. ordinary [434, 406, 260]. oriented [52, 281, 51]. orthogonal [298]. orthogonality [415]. oscillations [411]. oscillatory [309]. outer [107]. Output [204, 222, 323, 140]. overdetermined [56]. overlapping [222].

packet [51]. packet-oriented [51]. Padé [434]. PageRank [107]. parabolic [283, 132, 409]. paradigm [351]. parallel [164, 108, 242, 376]. Parameter [373, 163, 413, 183]. parameter-adjusting [183]. Parameter-related [373]. parameterization [89]. parameters [394, 65]. Parametric [13, 323, 101, 407, 22, 307, 24, 191]. parametrization [338]. Partial [141, 157, 234, 123, 228, 285, 225]. partially [157, 29]. particle [434, 413]. partitioning [122]. parts [382]. Pasciak [105]. Pasciak-like [105]. Passive [364]. Path [84, 394, 440, 166]. path-following [440]. pattern [410]. PDE [97, 345]. PDE-constrained [345]. peer [43, 44]. peer-to-peer [43, 44]. penalization [134]. penalized [278]. penalties [190]. penalty [192, 224]. Performance [79, 52, 266, 12, 43]. periodic [226, 258, 252, 245]. perishable [268]. Perron [125]. Perturbation [92, 21, 261, 294, 303, 194, 32]. perturbation-based [194]. perturbations [277]. perturbed [405, 393, 406, 226, 363]. phase [431, 292, 122, 272]. phenomenon [17]. photonic [51]. PID [385]. piecewise [181, 185]. pinning [261]. placement [328]. plankton [365]. planning [84]. plate [259]. players [374]. POD [237]. point [314, 199, 243, 20, 407, 299, 149, 253, 287, 184, 260, 333, 430, 105, 387, 254, 191, 57]. points [146, 165]. policy [282, 343, 425]. polyhedral [37]. polynomial [49, 241, 270, 379, 162]. polynomials [294, 161, 219]. population [319]. porous [336]. port [154]. Portfolio [66, 265, 317]. posedness [147]. Positive [305, 104, 6, 373]. positive-definite [104]. posteriori [85]. potential [128]. power [179, 261, 411, 290, 326, 364]. PR [50]. pre [425]. pre-sale [425]. Preconditioned [416, 297, 217, 337]. preconditioner [105]. predator [399]. predicting [362]. prediction [14]. predictor [314, 354]. predictor-corrector [314]. Preface [189, 288, 366, 121, 417, 180, 109, 1, 93, 23, 83, 320, 70, 40]. preinvex [401]. prescribed [29]. presence [249, 276]. preservation [306]. prey [399]. price [95]. Pricing [290]. Primal [199, 190, 407, 184, 36, 191]. Primal-dual [199, 190, 184, 36, 191]. principle [119]. priority [164, 47]. probabilistically [100, 94]. probabilities [49]. probability [29]. problem [394, 173, 271, 283, 295, 344, 223, 257, 357, 285, 317, 406, 247, 69, 311, 95, 37, 297, 373, 102, 187, 122, 128, 194, 437, 78, 228, 388, 349, 419, 268, 86, 235, 341, 368, 358, 64, 195]. problems [61, 414, 338, 88, 66, 298, 67, 100, 342, 392, 239, 25, 39, 179, 134, 274, 242, 118, 77, 273, 91, 275, 396, 94, 313, 427, 429, 116, 202, 204, 424, 166, 422, 245, 188, 89, 151, 416, 304, 430, 105, 241, 307, 224, 256, 34, 2, 415, 137, 81, 262, 205, 367, 345, 272, 85, 190, 208, 346, 68, 376, 176, 55, 309, 8, 57, 178, 152]. procedure [267]. process [174, 211]. processes [277, 226, 418]. processing [187]. processor [164]. product [269, 103, 317, 49, 159]. product-forms [49]. product-type [103]. production [423]. products [268]. profit [347]. programming [439, 265, 66, 314, 164, 135, 126, 9, 172, 154, 426, 26, 408, 36, 307, 262, 24, 208, 68, 129, 8]. programs [198, 30, 5, 192, 240, 300, 97, 156]. progress [429]. projected [297, 81, 367].

Projection [218, 25, 242, 237, 373, 430, 306]. Projection-based [218, 373, 306]. proof [216]. proper [220, 136, 215, 152]. Properties [401, 135, 200, 285, 5, 150, 176, 377]. proportional [433]. protected [365]. protocol [264]. prox [134]. prox-penalization [134]. Proximal [20, 188, 372, 299, 403]. proximity [165]. PRP [398]. pseudo [437]. pseudo-monotone [437]. pseudocontractive [174]. Pseudoconvexity [200]. pseudomonotonicity [299]. PSO [411]. public [420]. Pythagoras [219].

q [277]. q-uniformly [277]. QNMs [54]. QP [192]. QP-free [192]. QR [21]. Quadratic [87, 350, 414, 88, 199, 115, 198, 120, 243, 406, 419, 85, 197, 358, 8, 63]. quadrature [145]. quality [65]. Quantitative [300]. quantum [440, 114]. quasi [6, 353, 146, 201]. quasi- [146]. quasi-Newton [6, 353, 201]. quasiconvex [75]. quasiequilibrium [346]. Quasilinear [349, 289]. quaternionic [355]. queue [50]. queues [46].

Rachford [370]. radius [28]. rainfall [72]. randomized [370]. range [136, 182]. Rank [249, 193, 318, 305, 382, 205]. Rank-based [249]. Rank-one [193]. ranking [267]. ranks [382]. rate [79, 397, 330]. ratio [399]. ratio-dependent [399]. ratios [312]. reachable [144]. reaction [226, 263]. reaction-diffusion [263]. readers [79]. real [148, 154]. real-time [154]. reality [69, 116, 202, 204]. realization [433]. realizations [229]. reasoning [269]. Receptor [276]. recourse [198]. recovery [157]. recurrent [169]. recycling [423]. reduction [213, 257, 218, 316, 325, 234, 306, 85]. reductions [273]. Reeves [7].

Reeves-Type [7]. reference [140]. reflexive [18]. region [371, 181, 10, 129, 124, 196, 131, 201]. regional [409]. registration [86]. regression [280, 278]. regular [135]. Regularization [5]. regularized [390]. regulation [204, 212]. regulator [19]. reinsurance [214]. related [67, 74, 373, 28]. relating [24]. relation [296]. relations [97]. relaxation [197]. reliable [111]. rendezvous [319]. repair [295]. repositioning [420]. representation [331, 135, 307]. reprojection [315]. rescaling [20]. research [289, 377]. reservation [45]. resilient [389, 322]. resistor [171]. Resource [356, 347]. response [297, 399]. restoration [380]. restriction [217]. Results [245, 218, 258, 165, 418, 28, 89, 303, 33]. retrial [46]. retrospective [131]. revenue [281]. reversal [295]. review [428]. revisited [127]. RHUM [409]. Riccati [400, 237, 143]. rigid [321]. rigidity [321]. rings [162]. risk [389, 374]. risk-averse [374]. RLS [270]. road [45, 119]. robot [58, 160, 387]. robots [84]. Robust [230, 428, 168, 317, 424, 389, 268]. robustness [123, 163, 186]. rolling [387]. route [295]. routing [295, 268]. royal [119]. Runge [309].

S [361]. SAA [64]. Saddle [243, 105]. Saddle-point [243]. Safe [111]. sale [425]. saturated [397]. saturating [82]. saturation [222]. scalarization [175, 346]. scale [279, 237, 258, 296, 81, 360]. scales [118]. scaling [390]. scheduler [48]. scheduling [164, 187, 51]. scheme [336, 260, 240, 177]. Schilders [171]. Schulz [351]. search [14, 287, 381, 169, 137]. secant [296]. Second [313, 436, 406, 247, 240, 421, 177, 32, 55]. second-order [406, 421, 32, 55]. seepage [208]. segmentation [339]. Seidel [376].

SEIQR [206]. SEIQR-V [206]. seismically [395]. Selection [66, 383]. self [252, 384, 196]. self-adjoint [252]. self-consistent [384]. self-correcting [196]. Semi [154, 310, 63]. Semi-definite [154, 63]. Semi-local [310]. semicontinuity [101]. semidefinite [108, 135, 287, 373, 2, 191, 197, 68]. sensing [236]. Sensitivity [117, 342, 229]. sensor [328, 53]. separable [126, 129, 376]. sequence [243]. sequential [286]. series [362]. server [50]. service [52]. services [44]. set [356, 147, 427, 29, 190, 176, 215, 339, 178]. set-valued [147, 176, 215, 178]. sets [144, 307, 346]. setting [347]. setup [187]. shadowing [328]. shaped [27]. sharing [420]. sheet [333]. shop [271]. shortest [394]. shrinking [333]. signal [157, 167]. simplex [262]. simulated [72]. Simulation [99, 206]. simulations [387]. simultaneous [15]. Single [187, 50]. single-server [50]. Singular [329, 243, 87, 67, 261, 294, 386, 143, 378, 365]. Singularly [226, 393, 406, 363]. SISO [251]. sixth [57]. size [95]. skew [104]. skew-Hermitian [104]. Sliding [361]. slip [333]. small [413]. smart [259, 326]. smooth [277, 283, 181, 192, 307, 359]. smoothing [340, 37, 188, 224, 359, 55, 64]. snow [185]. snow-ice [185]. social [236]. soft [179, 247]. soft-computing [179]. software [275]. solution [233, 115, 330, 357, 305, 382, 406, 69, 250, 226, 430, 307, 86, 341]. Solutions [133, 88, 147, 73, 429, 424, 175, 128, 334, 345, 176]. Solvability [61]. solve [434, 414]. solver [98, 427]. solvers [351]. Solving [400, 239, 223, 311, 413, 208, 392, 25, 344, 179, 274, 242, 246, 294, 275, 316, 408, 304, 337, 403, 2, 262, 272, 106, 309]. Some [158, 74, 440, 429, 28, 150, 67, 145]. source [328, 42]. space [400, 393, 169, 177]. spaces [277, 31, 39, 344, 27, 146, 165, 174, 404, 17, 18]. Sparse [171, 323, 6, 286, 316, 157, 360, 193]. specific [331, 149]. specified [251]. spectral [28, 308, 296, 81, 367]. Speeding [195]. spline [125]. splines [142]. split [437]. splitting [104, 134, 274, 130]. splittings [220, 136]. SQP [56]. square [219]. squares [369, 373, 304, 368, 3]. SSOR [337]. Stability [22, 333, 306, 17, 198, 120, 300, 177, 346]. stabilizability [228]. Stabilization [354, 321, 16, 254]. stabilizer [137]. stage [198, 266, 399, 97, 374]. stagnation [333]. star [27]. star-shaped [27]. STATCOM [411]. state [277, 406, 427, 169, 361, 49, 210, 256]. state-dependent [277, 210]. state-linear [256]. stationary [127, 50]. statistical [48]. steady [49]. steady-state [49]. step [440, 145, 149, 253, 27, 296, 170]. stiff [248]. Stochastic [172, 265, 4, 198, 435, 429, 69, 116, 204, 240, 300, 408, 50, 97, 44, 211, 302, 64, 374]. stock [244]. stock-dependent [244]. strategies [226, 90, 170, 153]. strategy [291, 381, 385, 235, 420, 186]. streaming [44]. stretching [333]. stretching/shrinking [333]. Strict [18, 16]. strictly [284]. Strong [27, 174, 38, 134]. structural [232]. structure [67, 329, 14, 399, 251, 129, 376]. structured [213, 279, 129, 216]. structures [258]. study [47, 90, 270]. sub [295]. sub-route [295]. subdifferential [61]. submatrices [305]. submodular [190]. subset [24]. Subspace [124, 217]. substitution [126]. successive [68]. Sufficient [255, 77, 24, 162]. sum [243, 219]. super [43]. super-node [43]. supercomputer [275]. supply [96, 281, 312, 389, 266]. support [12]. surface [208]. Survey [375, 100, 429, 82, 304, 205]. survival [402]. sustainable [269, 389]. swarm [434, 413]. sweeping [277]. switched [354]. switches [51]. symmetric [148, 149, 166, 422, 7, 151, 337, 106, 309, 114]. symmetry [91]. symplectic [309]. synchronisation [261]. synthesis [251]. System [335, 431, 164, 147, 275, 141, 370, 423, 247, 340, 232, 185, 209, 241, 337, 210, 433, 163, 263, 212]. systems [323, 79, 19, 213, 279, 58, 103, 108, 329, 15, 435, 285, 393, 181, 229, 218, 409, 250, 348, 252, 234, 22, 231, 361, 363, 386, 82, 222, 436, 80, 112, 140, 364, 289, 324, 139, 354, 182, 420, 360, 106, 379, 17, 264, 378]. systems-homogenization [252].

T [361, 276]. T-S [361]. Tabu [14]. Tanh [225]. Tapia [8]. tardy [271]. target [280, 347]. target-environment [280]. targeted [30]. task [58, 286]. task-flow [286]. taxation [365]. Taylor [405]. technique [364, 359]. techniques [410, 258, 85]. temperature [332]. temperature-dependent [332]. temporarily [319]. tensor [400, 28, 367]. tensors [371, 284, 148]. term [191]. terminally [256]. terminals [154]. termination [59]. terms [372]. theorem [91, 146]. theorems [424, 422, 27]. theoretic [110]. theory [261, 329, 440, 142, 226, 428, 128, 119, 402]. therapy [276]. thermal [61, 185]. thermodynamic [185]. three [145, 226, 281, 27, 80, 263]. three-dimensional [226, 80, 263]. three-step [145, 27]. thresholding [318]. time [400, 118, 15, 91, 53, 275, 218, 154, 48, 383, 249, 69, 116, 204, 348, 234, 363, 362, 436, 123, 433, 268, 354]. time-access [48]. time-consuming [275]. time-delay [363]. time-driven [53]. time-invariant [436]. time-varying [218, 383, 436, 354]. times [187]. timetabling [173]. Todd [149]. Toeplitz [67]. Toeplitz-related [67]. Toland [156]. tolerant [378]. Topological [176]. topology [81]. total [41, 159, 368]. total-colorings [159]. trace [158]. tractor [154]. trade [282, 343]. trade-credit [282, 343]. trains [400]. trajectories [137]. trajectory [117]. transfer [383]. transferable [418]. transferable-utility [418]. transfers [33]. transform [76]. transit [295]. transmission [79]. transportation [79]. treatment [291, 90]. triality [128]. tridiagonal [238]. trigonometric [191]. trip [268]. truncated [354]. trust [371, 10, 129, 124, 196, 131, 201]. trust-region [10, 124]. tuberculosis [90]. Tug [395]. tuning [394]. turnpike [17]. twice [313]. Two [97, 170, 374, 271, 198, 6, 74, 336, 258, 292, 418, 155, 122, 244, 282, 350, 358, 8, 57, 264]. two-dimension [264]. two-dimensional [6]. two-echelon [244]. two-machine [271]. two-phase [292, 122]. two-point [57]. two-scale [258]. Two-stage [97, 374, 198]. Two-step [170]. Type [7, 314, 103, 336, 353, 362, 436, 306, 264].

Unbounded [277]. uncertain [361, 363, 386, 182, 378]. uncertainties [323]. uncertainty [424, 266, 374]. unconstrained [398, 148, 381, 296, 10, 124, 131]. uncontrolled [230]. underestimators [424]. Unified [346, 35, 125, 54]. uniformly [277, 174]. units [267]. Univariate [60]. univexity [26]. unknowns [336]. unrelated [164]. unstable [213]. updates [6]. updating [353]. upper [308]. urban [295]. useful [158]. using [269, 144, 279, 392, 400, 179, 65, 261, 286, 317, 424, 311, 160, 232, 203, 266, 333, 45, 413, 364, 334, 170]. utility [418].

V [206]. value [67, 340, 389, 421, 403, 349, 57]. valued [147, 205, 176, 215, 178]. values [261]. Variable [419, 372, 9, 406, 244, 282, 343]. variables [155, 266, 97]. variational [4, 147, 118, 91, 258, 252, 22, 437, 430, 335, 256, 216, 18, 374, 114]. variations [119]. varying [218, 383, 436, 354]. vector [31, 35, 313, 12, 345, 32, 346, 215, 152]. vehicle [268]. verification [351]. vertex [159]. vertical [300]. vertices [301]. via [158, 125, 77, 348, 381, 231, 386, 50, 59, 241, 346, 153]. viability [181]. vibration [259]. viral [397]. virtual [347]. viscosity [332]. visiting [41]. Volterra [405, 246, 334].

walking [13]. war [395]. water [65, 96, 11]. water-filling [11]. wavelet [336, 132, 311, 348]. wavelet-like [336]. wavelets [438]. Weak [134, 422, 296, 32]. wedge [196]. weed [317]. Weighted [301, 166, 234, 169, 418, 419]. weighted-path-following [166]. weights [356]. well [147, 128]. whale [383]. where [357]. white [211]. wild [230]. windows [268]. wireless [53, 42, 45]. wise [381]. without [192]. wolf [262]. work [46]. WVO [301].

Yang [133].

zero [243, 231]. zero-forcing [231]. zerosum [243]. zone [365]. zooming [170].

References

Kanzow:2011:P

[1] Christian Kanzow, Dong-Hui Li, and Nobuo Yamashita. Preface. Numerical Algebra, Control and Optimization, 1(1):i–v, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                             //aimsciences.org/article/doi/10.
                                             3934/naco.2011.1.1i.

Sun:2011:MSN

[2] Jie Sun. On methods for solving nonlinear semidefinite optimization problems. Numerical Algebra, Control and Optimization, 1(1):1–14, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.1.

Yuan:2011:RAN

[3] Ya-Xiang Yuan. Recent advances in numerical methods for nonlinear equations and nonlinear least squares. Numerical Algebra, Control and Optimization, 1(1):15–34, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.15.

Chen:2011:CBF

[4] Xiaojun Chen and Guihua Lin. CVaRbased formulation and approximation method for stochastic variational inequalities. Numerical Algebra, Control and Optimization, 1(1):35–48, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL

http://aimsciences.org/article/
doi/10.3934/naco.2011.1.35.

Hoheisel:2011:ICP

[5] Tim Hoheisel, Christian Kanzow, and Alexandra Schwartz. Improved convergence properties of the Lin–Fukushimaregularization method for mathematical programs with complementarity constraints. Numerical Algebra, Control and Optimization, 1(1):49–60, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.49.

Dai:2011:CAS

[6] Yuhong Dai and Nobuo Yamashita. Convergence analysis of sparse quasi-Newton updates with positive definite matrix completion for twodimensional functions. Numerical Algebra, Control and Optimization, 1 (1):61–69, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.61.

Li:2011:MFR

[7] Dong-Hui Li and Xiao-Lin Wang. A modified Fletcher–Reeves-type derivative-free method for symmetric nonlinear equations. Numerical Algebra, Control and Optimization, 1 (1):71–82, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.71.

Zhang:2011:CDT

[8] Ailing Zhang and Shunsuke Hayashi. Celis–Dennis–Tapia based approach to quadratic fractional programming problems with two quadratic constraints. Numerical Algebra, Control and Optimization, 1(1):83–98, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.83.

Hedar:2011:FBG

[9] Abdel-Rahman Hedar and Alaa Fahim. Filter-based genetic algorithm for mixed variable programming. Numerical Algebra, Control and Optimization, 1 (1):99–116, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.99.

Takaki:2011:DFT

[10] Jun Takaki and Nobuo Yamashita. A derivative-free trust-region algorithm for unconstrained optimization with controlled error. Numerical Algebra, Control and Optimization, 1 (1):117–145, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.117.

He:2011:NEI

[11] Simai He, Min Li, Shuzhong Zhang, and Zhi-Quan Luo. A nonconvergent example for the iterative water-filling algorithm. Numerical Algebra, Control and Optimization, 1(1):147–150, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.147.

Tatsumi:2011:PEM

[12] Keiji Tatsumi, Masashi Akao, Ryo Kawachi, and Tetsuzo Tanino. Performance evaluation of multiobjective multiclass support vector machines maximizing geometric margins. Numerical Algebra, Control and Optimization, 1(1):151–169, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.151.

Harata:2011:PEB

[13] Yuji Harata, Yoshihisa Banno, and Kouichi Taji. Parametric excitation based bipedal walking: Control method and optimization. Numerical Algebra, Control and Optimization, 1(1):171–190, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.171.

Hedar:2011:GAT

[14] Abdel-Rahman Hedar, Ahmed Fouad Ali, and Taysir Hassan Abdel-Hamid. Genetic algorithm and tabu search based methods for molecular 3D-structure prediction. Numerical Algebra, Control and Optimization, 1(1):191–209, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.191.

Fair:2011:AIF

[15] Martene L. Fair and Stephen L. Campbell. Active incipient fault detection in continuous time systems with multiple simultaneous faults. Numerical Algebra, Control and Optimization, 1(2):211–224, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.211.

Dick:2011:SLF

[16] Markus Dick, Martin Gugat, and Günter Leugering. A strict H^{1}-lyapunov function and feedback stabilization for the isothermal Euler equations with friction. Numerical Algebra, Control and Optimization, 1(2):225–244, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.225.

Zaslavski:2011:STP

[17] Alexander J. Zaslavski. Stability of a turnpike phenomenon for a class of optimal control systems in metric spaces. Numerical Algebra, Control and Optimization, 1(2):245–260, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.245.

Zhong:2011:SFG

[18] Ren-You Zhong and Nan-Jing Huang. Strict feasibility for generalized mixed variational inequality in reflexive Banach spaces. Numerical Algebra, Control and Optimization, 1(2):261–274, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL

http://aimsciences.org/article/
doi/10.3934/naco.2011.1.261.

Basin:2011:OIC

[19] Michael Basin and Pablo Rodriguez-Ramirez. An optimal impulsive control regulator for linear systems. Numerical Algebra, Control and Optimization, 1(2):275–282, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.275.

Griva:2011:PPN

[20] Igor Griva and Roman A. Polyak. Proximal point nonlinear rescaling method for convex optimization. Numerical Algebra, Control and Optimization, 1(2):283–299, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.283.

Chang:2011:MPA

[21] Xiao-Wen Chang and Ren-Cang Li. Multiplicative perturbation analysis for QR factorizations. Numerical Algebra, Control and Optimization, 1 (2):301–316, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.301.

Li:2011:SAP

[22] Shengji Li, Chunmei Liao, and Minghua Li. Stability analysis of parametric variational systems. Numerical Algebra, Control and Optimization, 1(2):317–331, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.317.

Li:2011:P

[23] Shengji Li, Nan-Jing Huang, and Xinmin Yang. Preface. Numerical Algebra, Control and Optimization, 1(3):i–ii, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.3i.

Verma:2011:GPS

[24] Ram U. Verma. General parametric sufficient optimality conditions for multiple objective fractional subset programming relating to generalized (ρ, η, A)-invexity. Numerical Algebra, Control and Optimization, 1(3):333–339, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.333.

Ceng:2011:EPM

[25] Luchuan Ceng, Qamrul Hasan Ansari, and Jen-Chih Yao. Extragradientprojection method for solving constrained convex minimization problems. Numerical Algebra, Control and Optimization, 1(3):341–359, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.341.

Long:2011:OCD

[26] Xian-Jun Long and Jing Quan. Optimality conditions and duality for minimax fractional programming involving nonsmooth generalized univexity. Numerical Algebra, Control and Optimization, 1(3):361–370, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.361.

Lee:2011:SCT

[27] Byung-Soo Lee. Strong convergence theorems with three-step iteration in star-shaped metric spaces. Numerical Algebra, Control and Optimization, 1(3):371–379, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.371.

Ling:2011:SRE

[28] Chen Ling and Liqun Qi. Some results on l^{k}-eigenvalues of tensor and related spectral radius. Numerical Algebra, Control and Optimization, 1(3):381–388, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.381.

Studniarski:2011:FAM

[29] Marcin Studniarski. Finding all minimal elements of a finite partially ordered set by genetic algorithm with a prescribed probability. Numerical Algebra, Control and Optimization, 1(3):389–398, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.389.

Ding:2011:ITC

[30] Mingfang Ding, Yanqun Liu, and John Anthony Gear. An improved targeted climbing algorithm for linear programs. Numerical Algebra, Control and Optimization, 1(3):399–405, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.399.

Bot:2011:LVO

[31] Radu Ioan Boţ and Sorin-Mihai Grad. On linear vector optimization duality in infinite-dimensional spaces. Numerical Algebra, Control and Optimization, 1(3):407–415, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.407.

Wang:2011:CSO

[32] Qilin Wang, Shengji Li, and Kok Lay Teo. Continuity of second-order adjacent derivatives for weak perturbation maps in vector optimization. Numerical Algebra, Control and Optimization, 1(3):417–433, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.417.

Pontani:2011:OTO

[33] Mauro Pontani. Orbital transfers: optimization methods and recent results. Numerical Algebra, Control and Optimization, 1(3):435–485, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.435.

Stipanovic:2011:NMA

[34] Dušan M. Stipanović, Claire J. Tomlin, and George Leitmann. A note on monotone approximations of minimum and maximum functions and multi-objective problems. Numerical Algebra, Control and Optimization, 1(3):487–493, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.487.

Chen:2011:UNA

[35] Chunrong Chen. A unified nonlinear augmented Lagrangian approach for nonconvex vector optimization. Numerical Algebra, Control and Optimization, 1(3):495–508, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.495.

Pillo:2011:PDA

[36] Gianni Di Pillo, Giampaolo Liuzzi, and Stefano Lucidi. A primal-dual algorithm for nonlinear programming exploiting negative curvature directions.

Numerical Algebra, Control and Optimization, 1(3):509–528, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.509.

Li:2011:SBL

[37] Yafeng Li, Guo Sun, and Yiju Wang. A smoothing Broyden-like method for polyhedral cone constrained eigenvalue problem. Numerical Algebra, Control and Optimization, 1(3): 529–537, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.529.

Burachik:2011:ASD

[38] Regina S. Burachik and Xiaoqi Yang. Asymptotic strong duality. Numerical Algebra, Control and Optimization, 1(3):539–548, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.539.

Chadli:2011:BME

[39] Ouayl Chadli, Hicham Mahdioui, and Jen-Chih Yao. Bilevel mixed equilibrium problems in Banach spaces: existence and algorithmic aspects. Numerical Algebra, Control and Optimization, 1(3):549–561, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.549.

Yue:2011:P

[40] Wuyi Yue, Herwig Bruneel, Bong Dae Choi, and Shoji Kasahara. Preface.

Numerical Algebra, Control and Optimization, 1(4):i–ii, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.4i.

Kabir:2011:OVO

[41] K. Habibul Kabir, Masahiro Sasabe, and Tetsuya Takine. Optimal visiting order of isolated clusters in DTNs to minimize the total mean delivery delay of bundles. Numerical Algebra, Control and Optimization, 1(4):563–576, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.563.

Minami:2011:AMS

[42] Keisuke Minami, Takahiro Matsuda, Tetsuya Takine, and Taku Noguchi. Asynchronous multiple source network coding for wireless broadcasting. Numerical Algebra, Control and Optimization, 1(4):577–592, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.577.

Kuraya:2011:LDP

[43] Kazuhiko Kuraya, Hiroyuki Masuyama, and Shoji Kasahara. Load distribution performance of super-node based peerto-peer communication networks: A nonstationary Markov chain approach. Numerical Algebra, Control and Optimization, 1(4):593–610, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.593.

Senda:2011:SFM

[44] Shuichiro Senda, Hiroyuki Masuyama, and Shoji Kasahara. A stochastic fluid model for on-demand peerto-peer streaming services. Numerical Algebra, Control and Optimization, 1(4):611–626, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.611.

Park:2011:ACD

[45] Jin Soo Park, Kyung Jae Kim, Yun Han Bae, and Bong Dae Choi. Admission control by dynamic bandwidth reservation using road layout and bidirectional navigator in wireless multimedia networks. Numerical Algebra, Control and Optimization, 1(4):627–638, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.627.

Phung-Duc:2011:MRQ

[46] Tuan Phung-Duc and Ken’ichi Kawanishi. Multiserver retrial queues with after-call work. Numerical Algebra, Control and Optimization, 1(4): 639–656, ???? 2011. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2011.1.639.

Maertens:2011:CDD

[47] Tom Maertens, Joris Walraevens, and Herwig Bruneel. Controlling delay differentiation with priority jumps: Analytical study. Numerical Algebra, Control and Optimization, 1(4):

657–673, ????  2011.  CODEN

???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.657.

Ishizaki:2011:AST

[48] Fumio Ishizaki. Analysis of the statistical time-access fairness index of one-bit feedback fair scheduler. Numerical Algebra, Control and Optimization, 1(4):675–689, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.675.

Liu:2011:KPF

[49] Hsin-Yi Liu and Hsing Paul Luh. Kronecker product-forms of steady-state probabilities with C_{k}/C_{m}/1 by matrix polynomial approaches. Numerical Algebra, Control and Optimization, 1(4):691–711, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.691.

Miyoshi:2011:SLP

[50] Naoto Miyoshi. On the stationary LCFS-PR single-server queue: A characterization via stochastic intensity. Numerical Algebra, Control and Optimization, 1(4):713–725, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.713.

Rogiest:2011:OPO

[51] Wouter Rogiest, Koen De Turck, Koenraad Laevens, Dieter Fiems, Sabine Wittevrongel, and Herwig Bruneel. On the optimality of packet-oriented scheduling in photonic switches with delay lines. Numerical Algebra, Control and Optimization, 1(4):727–747, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.727.

Jin:2011:PEC

[52] Shunfu Jin, Wuyi Yue, and Zhanqiang Huo. Performance evaluation for connection oriented service in the next generation Internet. Numerical Algebra, Control and Optimization, 1(4):749–761, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.749.

Galmes:2011:MCN

[53] Sebastià Galmés. Markovian characterization of node lifetime in a timedriven wireless sensor network. Numerical Algebra, Control and Optimization, 1(4):763–780, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.763.

Kouvatsos:2011:UMA

[54] Demetres D. Kouvatsos, Jumma S. Alanazi, and Kevin Smith. A unified ME algorithm for arbitrary open QNMs with mixed blocking mechanisms. Numerical Algebra, Control and Optimization, 1(4):781–816, ???? 2011. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2011.1.781.

Yuan:2012:SNM

[55] Yanhong Yuan, Hongwei Zhang, and Liwei Zhang. A smoothing Newton method for generalized Nash equilibrium problems with second-order cone constraints. Numerical Algebra, Control and Optimization, 2(1):1–18, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.1.

Hao:2012:GCS

[56] Chunlin Hao and Xinwei Liu. Global convergence of an SQP algorithm for nonlinear optimization with overdetermined constraints. Numerical Algebra, Control and Optimization, 2 (1):19–29, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.19.

Zhang:2012:SON

[57] Xiao-Yu Zhang and Qing Fang. A sixth order numerical method for a class of nonlinear two-point boundary value problems. Numerical Algebra, Control and Optimization, 2 (1):31–43, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.31.

Chan:2012:ABO

[58] Raymond Ching Man Chan and Henry Ying Kei Lau. An AIS-based optimal control framework for longevity and task achievement of multi-robot systems. Numerical Algebra, Control and Optimization, 2(1):45–56, ???? 2012.

CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.45.

Ong:2012:GOD

[59] Bun Theang Ong and Masao Fukushima. Global optimization via differential evolution with automatic termination. Numerical Algebra, Control and Optimization, 2(1):57–67, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.57.

Kvasov:2012:UGL

[60] Dmitri E. Kvasov and Yaroslav D. Sergeyev. Univariate geometric Lipschitz global optimization algorithms. Numerical Algebra, Control and Optimization, 2(1):69–90, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.69.

Adly:2012:SCT

[61] Samir Adly, Oanh Chau, and Mohamed Rochdi. Solvability of a class of thermal dynamical contact problems with subdifferential conditions. Numerical Algebra, Control and Optimization, 2(1):91–104, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.91.

Albrecht:2012:BOA

[62] Sebastian Albrecht, Marion Leibold, and Michael Ulbrich. A bilevel optimization approach to obtain optimal cost functions for human arm movements. Numerical Algebra, Control and Optimization, 2(1):105–127, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.105.

Zhang:2012:EAC

[63] Lipu Zhang, Yinghong Xu, and Zhengjing Jin. An efficient algorithm for convex quadratic semi-definite optimization. Numerical Algebra, Control and Optimization, 2(1):129–144, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.129.

Zhang:2012:CSS

[64] Jie Zhang, Yue Wu, and Liwei Zhang. A class of smoothing SAA methods for a stochastic linear complementarity problem. Numerical Algebra, Control and Optimization, 2(1):145–156, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.145.

Chen:2012:IWQ

[65] Guangzhou Chen, Guijian Liu, Jiaquan Wang, and Ruzhong Li. Identification of water quality model parameters using artificial bee colony algorithm. Numerical Algebra, Control and Optimization, 2(1):157–165, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.157.

An:2012:DPA

[66] Le Thi Hoai An, Tran Duc Quynh, and Pham Dinh Tao. A DC programming approach for a class of bilevel programming problems and its application in portfolio selection. Numerical Algebra, Control and Optimization, 2(1):167–185, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.167.

Bai:2012:SIS

[67] Zheng-Jian Bai, Xiao-Qing Jin, and Seak-Weng Vong. On some inverse singular value problems with Toeplitz-related structure. Numerical Algebra, Control and Optimization, 2(1):187–192, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.187.

Xu:2012:FSL

[68] Yi Xu and Wenyu Sun. A filter successive linear programming method for nonlinear semidefinite programming problems. Numerical Algebra, Control and Optimization, 2(1): 193–206, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.193.

Kek:2012:FSN

[69] Sie Long Kek, Kok Lay Teo, and Mohd Ismail Abd Aziz. Filtering solution of nonlinear stochastic optimal control problem in discrete-time with model-reality differences. Numerical Algebra, Control and Optimization, 2(1):207–222, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.207.

Wang:2012:P

[70] Song Wang and Yong Hong Wu. Preface. Numerical Algebra, Control and Optimization, 2(2):i–ii, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.2i.

Cizmesija:2012:FMG

[71] Aleksandra Čižmešija, Iva Franjić, Josip Pečarić, and Dora Pokaz. On a family of means generated by the Hardy– Littlewood maximal inequality. Numerical Algebra, Control and Optimization, 2(2):223–231, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.223.

Piantadosi:2012:MEM

[72] Julia Piantadosi, Phil Howlett, Jonathan Borwein, and John Henstridge. Maximum entropy methods for generating simulated rainfall. Numerical Algebra, Control and Optimization, 2(2):233–256, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.233.

Cohen:2012:MSM

[73] Samuel N. Cohen and Lukasz Szpruch. On Markovian solutions to Markov Chain BSDEs. Numerical Algebra, Control and Optimization, 2(2): 257–269, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.257.

Dragomir:2012:SNB

[74] S. S. Dragomir and I. Gomm. Some new bounds for two mappings related to the Hermite–Hadamard inequality for convex functions. Numerical Algebra, Control and Optimization, 2(2):271–278, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.271.

Dragomir:2012:JIQ

[75] S. S. Dragomir and C. E. M. Pearce. Jensen’s inequality for quasiconvex functions. Numerical Algebra, Control and Optimization, 2(2):279–291, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.279.

Ejov:2012:HTM

[76] Vladimir Ejov and Anatoli Torokhti. How to transform matrices U_{1}, …, U_{p} to matrices V_{1}, …, V_{p} so that V_{i}V_{j}^{T} = O if i ≠ j ? Numerical Algebra, Control and Optimization, 2(2):293–299, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.293.

Eberhard:2012:SOC

[77] A. C. Eberhard and C. E. M. Pearce. A sufficient optimality condition for nonregular problems via a nonlinear Lagrangian. Numerical Algebra, Control and Optimization, 2(2):301–331, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.301.

Pearce:2012:DPM

[78] Charles E. M. Pearce, Krzysztof Szajowski, and Mitsushi Tamaki. Duration problem with multiple exchanges. Numerical Algebra, Control and Optimization, 2(2):333–355, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.333.

Baek:2012:PAT

[79] Sangkyu Baek, Jinsoo Park, and Bong Dae Choi. Performance analysis of transmission rate control algorithm from readers to a middleware in intelligent transportation systems. Numerical Algebra, Control and Optimization, 2(2):357–375, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.357.

Mohd–Mahali:2012:NME

[80] Shalela Mohd-Mahali, Song Wang, Xia Lou, and Sungging Pintowantoro. Numerical methods for estimating effective diffusion coefficients of threedimensional drug delivery systems. Numerical Algebra, Control and Optimization, 2(2):377–393, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.377.

Tavakoli:2012:NSP

[81] Rouhollah Tavakoli and Hongchao Zhang. A nonmonotone spectral projected gradient method for large-scale topology optimization problems. Numerical Algebra, Control and Optimization, 2(2):395–412, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.395.

Mahmoud:2012:CDL

[82] Magdi S. Mahmoud and Mohammed M. Hussain. Control design of linear systems with saturating actuators: A survey. Numerical Algebra, Control and Optimization, 2(2):413–435, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.413.

Rehbock:2012:P

[83] Volker Rehbock and Ryan Loxton. Preface. Numerical Algebra, Control and Optimization, 2(3):i, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.3i.

Gerdts:2012:PPC

[84] Matthias Gerdts, René Henrion, Dietmar Hömberg, and Chantal Landry. Path planning and collision avoidance for robots. Numerical Algebra, Control and Optimization, 2(3): 437–463, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.437.

Vossen:2012:MRT

[85] Georg Vossen and Stefan Volkwein. Model reduction techniques with aposteriori error analysis for linearquadratic optimal control problems. Numerical Algebra, Control and Optimization, 2(3):465–485, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.465.

Wagner:2012:DMS

[86] Marcus Wagner. A direct method for the solution of an optimal control problem arising from image registration. Numerical Algebra, Control and Optimization, 2(3):487–510, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.487.

Aronna:2012:QOC

[87] M. Soledad Aronna, J. Frédréic Bonnans, Andrei V. Dmitruk, and Pablo A. Lotito. Quadratic order conditions for bang-singular extremals. Numerical Algebra, Control and Optimization, 2(3):511–546, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.511.

Alt:2012:EBE

[88] Walter Alt, Robert Baier, Matthias Gerdts, and Frank Lempio. Error bounds for Euler approximation of linear-quadratic control problems with bang-bang solutions. Numerical Algebra, Control and Optimization, 2(3):547–570, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.547.

Loxton:2012:CPO

[89] Ryan Loxton, Qun Lin, Volker Rehbock, and Kok Lay Teo. Control parameterization for optimal control problems with continuous inequality constraints: New convergence results. Numerical Algebra, Control and Optimization, 2(3):571–599, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.571.

Silva:2012:OCS

[90] Cristiana J. Silva and Delfim F. M. Torres. Optimal control strategies for tuberculosis treatment: A case study in Angola. Numerical Algebra, Control and Optimization, 2(3):601–617, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.601.

Frederico:2012:NST

[91] Gastão S. F. Frederico and Delfim F. M. Torres. Noether’s symmetry theorem for variational and optimal control problems with time delay. Numerical Algebra, Control and Optimization, 2(3):619–630, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.619.

Schattler:2012:PFC

[92] Heinz Schättler and Urszula Ledzewicz. Perturbation feedback control: A geometric interpretation. Numerical Algebra, Control and Optimization, 2 (3):631–654, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.631.

Koch:2012:P

[93] Thorsten Koch and Xiaoling Sun. Preface. Numerical Algebra, Control and Optimization, 2(4):i–ii, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.4i.

Henrion:2012:GEG

[94] René Henrion. Gradient estimates for Gaussian distribution functions: application to probabilistically constrained optimization problems. Numerical Algebra, Control and Optimization, 2(4):655–668, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.655.

Kiessling:2012:ISP

[95] Miriam Kiessling, Sascha Kurz, and Jörg Rambau. The integrated size and price optimization problem. Numerical Algebra, Control and Optimization, 2(4):669–693, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.669.

Gleixner:2012:TGO

[96] Ambros M. Gleixner, Harald Held, Wei Huang, and Stefan Vigerske. Towards globally optimal operation of water supply networks. Numerical Algebra, Control and Optimization, 2(4):695–711, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.695.

Schultz:2012:TSS

[97] Rüdiger Schultz. Two-stage stochastic programs: Integer variables, dominance relations and PDE constraints. Numerical Algebra, Control and Optimization, 2(4):713–738, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.713.

Berthold:2012:ACI

[98] Timo Berthold, Ambros M. Gleixner, Stefan Heinz, and Stefan Vigerske. Analyzing the computational impact of MIQCP solver components. Numerical Algebra, Control and Optimization, 2(4):739–748, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.739.

Chen:2012:SLD

[99] Rachel Chen, Jianqiang Hu, and Yijie Peng. Simulation of Lévy–driven models and its application in finance. Numerical Algebra, Control and Optimization, 2(4):749–765, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.749.

Bai:2012:SPC

[100] Xiaodi Bai, Xiaojin Zheng, and Xiaoling Sun. A survey on probabilistically constrained optimization problems. Numerical Algebra, Control and Optimization, 2(4):767–778, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.767.

Chen:2012:NSP

[101] Chunrong Chen and Zhimiao Fang. A note on semicontinuity to a parametric generalized Ky Fan inequality. Numerical Algebra, Control and Optimization, 2(4):779–784, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.779.

Liu:2012:MMM

[102] Xin-Guo Liu and Kun Wang. A multigrid method for the maximal correlation problem. Numerical Algebra, Control and Optimization, 2 (4):785–796, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.785.

Chen:2012:PTG

[103] Fang Chen, Ning Gao, and Yao-Lin Jiang. On product-type generalized block AOR method for augmented linear systems. Numerical Algebra, Control and Optimization, 2(4):797–809, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.797.

Cao:2012:GPD

[104] Yang Cao, Wei-Wei Tan, and Mei-Qun Jiang. A generalization of the positivedefinite and skew-Hermitian splitting iteration. Numerical Algebra, Control and Optimization, 2(4):811–821, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.811.

Peng:2012:NBP

[105] Xiao-Fei Peng and Wen Li. A new Bramble–Pasciak-like preconditioner for saddle point problems. Numerical Algebra, Control and Optimization, 2(4):823–838, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.823.

Yang:2012:NMM

[106] Ai-Li Yang and Yu-Jiang Wu. Newton-MHSS methods for solving systems of nonlinear equations with complex symmetric Jacobian matrices. Numerical Algebra, Control and Optimization, 2(4):839–853, ???? 2012. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2012.2.839.

Bai:2012:CIO

[107] Zhong-Zhi Bai. On convergence of the inner–outer iteration method for computing PageRank. Numerical Algebra, Control and Optimization, 2 (4):855–862, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.855.

Cui:2012:CGP

[108] Yanxing Cui, Chuanlong Wang, and Ruiping Wen. On the convergence of generalized parallel multisplitting iterative methods for semidefinite linear systems. Numerical Algebra, Control and Optimization, 2(4):863–873, ???? 2012. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2012.2.863.

Goh:2013:P

[109] B. S. Goh.  Preface.  Numeri-

cal Algebra, Control and Optimization, 3(1):i–iii, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.1i.

Bhattacharya:2013:JMN

[110] Sourabh Bhattacharya, Abhishek Gupta, and Tamer Başar. Jamming in mobile networks: A game-theoretic approach. Numerical Algebra, Control and Optimization, 3(1):1–30, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.1.

Stipanovic:2013:SRC

[111] Dušan M. Stipanović, Christopher Valicka, Claire J. Tomlin, and Thomas R.

Bewley. Safe and reliable coverage control. Numerical Algebra, Control and Optimization, 3(1):31–48, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.31.

Morzfeld:2013:CDL

[112] Matthias Morzfeld, Daniel T. Kawano, and Fai Ma. Characterization of damped linear dynamical systems in free motion. Numerical Algebra, Control and Optimization, 3(1):49–62, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.49.

Wirl:2013:IGD

[113] Franz Wirl and Andreas J. Novak. Instability and growth due to adjustment costs. Numerical Algebra, Control and Optimization, 3(1):63–76, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.63.

daCruz:2013:HSQ

[114] Artur M. C. Brito da Cruz, Natália Martins, and Delfim F. M. Torres. Hahn’s symmetric quantum variational calculus. Numerical Algebra, Control and Optimization, 3(1):77–94, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.77.

Bensoussan:2013:LQD

[115] Alain Bensoussan, Shaokuan Chen, and Suresh P. Sethi. Linear quadratic differential games with mixed leadership: The open-loop solution. Numerical Algebra, Control and Optimization, 3(1):95–108, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.95.

Kek:2013:IAB

[116] Sie Long Kek, Mohd Ismail Abd Aziz, Kok Lay Teo, and Rohanin Ahmad. An iterative algorithm based on modelreality differences for discrete-time nonlinear stochastic optimal control problems. Numerical Algebra, Control and Optimization, 3(1):109–125, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.109.

McDonald:2013:SBT

[117] Dale McDonald. Sensitivity based trajectory following control damping methods. Numerical Algebra, Control and Optimization, 3(1):127–143, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.127.

Dryl:2013:NOC

[118] Monika Dryl and Delfim F. M. Torres. Necessary optimality conditions for infinite horizon variational problems on time scales. Numerical Algebra, Control and Optimization, 3(1):145–160, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.145.

Pesch:2013:CRR

[119] Hans Josef Pesch. Carathéodory’s royal road of the calculus of variations: Missed exits to the maximum principle of optimal control theory. Numerical Algebra, Control and Optimization, 3(1):161–173, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.161.

DAlto:2013:IQS

[120] Luis D’Alto and Martin Corless. Incremental quadratic stability. Numerical Algebra, Control and Optimization, 3(1):175–201, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.175.

Ding:2013:P

[121] Jiu Ding, Bingsheng He, Qin Ni, and Wenyu Sun. Preface. Numerical Algebra, Control and Optimization, 3(2):i–ii, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.2i.

Ma:2013:TPM

[122] Feng Ma and Mingfang Ni. A twophase method for multidimensional number partitioning problem. Numerical Algebra, Control and Optimization, 3(2):203–206, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.203.

Mao:2013:PEA

[123] Xiaobin Mao and Hua Dai. Partial eigenvalue assignment with time delay robustness. Numerical Algebra, Control and Optimization, 3(2): 207–221, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.207.

Zhang:2013:STR

[124] Xin Zhang, Jie Wen, and Qin Ni. Subspace trust-region algorithm with conic model for unconstrained optimization. Numerical Algebra, Control and Optimization, 3(2):223–234, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.223.

Ding:2013:UME

[125] Jiu Ding and Noah H. Rhee. A unified maximum entropy method via spline functions for Frobenius–Perron operators. Numerical Algebra, Control and Optimization, 3(2):235–245, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.235.

He:2013:LAD

[126] Bingsheng He and Xiaoming Yuan. Linearized alternating direction method of multipliers with Gaussian back substitution for separable convex programming. Numerical Algebra, Control and Optimization, 3(2):247–260, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.247.

Chen:2013:SIR

[127] Xuzhou Chen, Xinghua Shi, and Yimin Wei. The stationary iterations revisited. Numerical Algebra, Control and Optimization, 3(2):261–270, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.261.

Morales-Silva:2013:CST

[128] Daniel Morales-Silva and David Yang Gao. Complete solutions and triality theory to a nonconvex optimization problem with double-well potential in R^{n}. Numerical Algebra, Control and Optimization, 3(2):271–282, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.271.

Xue:2013:STR

[129] Dan Xue, Wenyu Sun, and Hongjin He. A structured trust region method for nonconvex programming with separable structure. Numerical Algebra, Control and Optimization, 3(2): 283–293, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.283.

Ding:2013:MFB

[130] Xiao Ding and Deren Han. A modification of the forward-backward splitting method for maximal monotone mappings. Numerical Algebra, Control and Optimization, 3(2):295–307, ???? 2013.

CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.295.

Zhao:2013:NRC

[131] Lijuan Zhao and Wenyu Sun. Nonmonotone retrospective conic trust region method for unconstrained optimization. Numerical Algebra, Control and Optimization, 3(2):309–325, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.309.

Guo:2013:AWM

[132] Qiang Guo and Dong Liang. An adaptive wavelet method and its analysis for parabolic equations. Numerical Algebra, Control and Optimization, 3(2):327–345, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.327.

Cibotarica:2013:SYB

[133] A. Cibotarica, Jiu Ding, J. Kolibal, and Noah H. Rhee. Solutions of the Yang– Baxter matrix equation for an idempotent. Numerical Algebra, Control and Optimization, 3(2):347–352, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.347.

Chbani:2013:WSC

[134] Zaki Chbani and Hassan Riahi. Weak and strong convergence of proxpenalization and splitting algorithms for bilevel equilibrium problems. Numerical Algebra, Control and Optimization, 3(2):353–366, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.353.

Dobre:2013:MPR

[135] Cristian Dobre. Mathematical properties of the regular *-representation of matrix ∗-algebras with applications to semidefinite programming. Numerical Algebra, Control and Optimization, 3(2):367–378, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.367.

Jena:2013:IRM

[136] Litismita Jena and Sabyasachi Pani. Index-range monotonicity and indexproper splittings of matrices. Numerical Algebra, Control and Optimization, 3(2):379–388, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.379.

Tarasyev:2013:ANS

[137] Alexander Tarasyev and Anastasia Usova. Application of a nonlinear stabilizer for localizing search of optimal trajectories in control problems with infinite horizon. Numerical Algebra, Control and Optimization, 3(3):389–406, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.389.

Pineda:2013:MCC

[138] M. Delgado Pineda, E. A. Galperin, and P. Jiménez Guerra. MAPLE code of the cubic algorithm for multiobjective optimization with box constraints. Numerical Algebra, Control and Optimization, 3(3):407–424, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.407.

Udwadia:2013:GNC

[139] Firdaus E. Udwadia and Thanapat Wanichanon. On general nonlinear constrained mechanical systems. Numerical Algebra, Control and Optimization, 3(3):425–443, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.425.

Nelson:2013:DMR

[140] James P. Nelson and Mark J. Balas.

Direct model reference adaptive control of linear systems with input/output delays. Numerical Algebra, Control and Optimization, 3(3):445–462, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.445.

Goh:2013:PNM

[141] B. S. Goh, W. J. Leong, and Z. Siri.

Partial Newton methods for a system of equations. Numerical Algebra, Control and Optimization, 3 (3):463–469, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.463.

Isaev:2013:ITS

[142] Vyacheslav K. Isaev and Vyacheslav V. Zolotukhin. Introduction to the theory of splines with an optimal mesh. Linear Chebyshev splines and applications. Numerical Algebra, Control and Optimization, 3(3):471–489, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.471.

Wang:2013:DIS

[143] Wei guo Wang, Wei chao Wang, and Ren cang Li. Deflating irreducible singular M-matrix algebraic Riccati equations. Numerical Algebra, Control and Optimization, 3(3):491–518, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.491.

Baier:2013:ARS

[144] Robert Baier, Matthias Gerdts, and Ilaria Xausa. Approximation of reachable sets using optimal control algorithms. Numerical Algebra, Control and Optimization, 3(3):519–548, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.519.

Khattri:2013:ANS

[145] Sanjay Khattri. Another note on some quadrature based three-step iterative methods for non-linear equations. Numerical Algebra, Control and Optimization, 3(3):549–555, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.549.

Lee:2013:CTC

[146] Byung-Soo Lee. A convergence theorem of common fixed points of a countably infinite family of asymptotically quasif_{i}-expansive mappings in convex metric spaces. Numerical Algebra, Control and Optimization, 3(3):557–565, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.557.

Chen:2013:ESW

[147] Jiawei Chen, Zhongping Wan, and Liuyang Yuan. Existence of solutions and α-well-posedness for a system of constrained set-valued variational inequalities. Numerical Algebra, Control and Optimization, 3(3):567–581, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.567.

Han:2013:UOA

[148] Lixing Han. An unconstrained optimization approach for finding real eigenvalues of even order symmetric tensors. Numerical Algebra, Control and Optimization, 3(3):583–599, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.583.

Kheirfam:2013:FNT

[149] Behrouz Kheirfam. A full Nesterov– Todd step infeasible interior-point algorithm for symmetric optimization based on a specific kernel function. Numerical Algebra, Control and Optimization, 3(4):601–614, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.601.

Yan:2013:SPC

[150] Lijia Yan. Some properties of a class of (F, E)-G generalized convex functions. Numerical Algebra, Control and Optimization, 3(4):615–625, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.615.

Miao:2013:EBS

[151] Xin-He Miao and Jein-Shan Chen. Error bounds for symmetric cone complementarity problems. Numerical Algebra, Control and Optimization, 3(4):627–641, ???? 2013. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.627.

Zhao:2013:CBP

[152] Kequan Zhao and Xinmin Yang. Characterizations of the E-Benson proper efficiency in vector optimization problems. Numerical Algebra, Control and Optimization, 3(4):643–653, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2013.3.643.

Xu:2013:DBE

[153] Honglei Xu, Peng Sui, Guanglu Zhou, and Louis Caccetta. Dampening bullwhip effect of order-up-to inventory strategies via an optimal control method. Numerical Algebra, Control and Optimization, 3 (4):655–664, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.655.

Huang:2013:SDP

[154] Wei Huang, Ka-Fai Cedric Yiu, and Henry Y. K. Lau. Semi-definite programming based approaches for realtime tractor localization in port container terminals. Numerical Algebra, Control and Optimization, 3(4): 665–680, ???? 2013. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2013.3.665.

Lyu:2014:HHI

[155] Shu-Lin Lyu. On the Hermite– Hadamard inequality for convex functions of two variables. Numerical Algebra, Control and Optimization, 4(1):1–8, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.1.

Zhou:2014:TFL

[156] Yuying Zhou and Gang Li. The Toland– Fenchel–Lagrange duality of DC programs for composite convex functions. Numerical Algebra, Control and Optimization, 4(1):9–23, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.9.

Kong:2014:PGP

[157] Lingchen Kong, Naihua Xiu, and Guokai Liu. Partial S-goodness for partially sparse signal recovery. Numerical Algebra, Control and Optimization, 4(1):25–38, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.25.

Chang:2014:SUI

[158] Yu-Lin Chang and Chin-Yu Yang. Some useful inequalities via trace function method in Euclidean Jordan algebras. Numerical Algebra, Control and Optimization, 4(1):39–48, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.39.

Tian:2014:AVD

[159] Shuangliang Tian, Ping Chen, Yabin Shao, and Qian Wang. Adjacent vertex distinguishing edge-colorings and totalcolorings of the Cartesian product of graphs. Numerical Algebra, Control and Optimization, 4(1):49–58, ???? 2014. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.49.

Ko:2014:GFB

[160] Chun-Hsu Ko and Jing-Kun Chen. Grasping force based manipulation for multifingered hand-arm robot using neural networks. Numerical Algebra, Control and Optimization, 4 (1):59–74, ???? 2014. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.59.

Ibrahim:2014:ALE

[161] Nur Fadhilah Ibrahim. An algorithm for the largest eigenvalue of nonhomogeneous nonnegative polynomials. Numerical Algebra, Control and Optimization, 4(1):75–91, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.75.

Zhang:2014:SCE

[162] Shenggui Zhang. A sufficient condition of Euclidean rings given by polynomial optimization over a box. Numerical Algebra, Control and Optimization, 4(2):93–101, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.93.

Wang:2014:PIN

[163] Lei Wang, Jinlong Yuan, Yingfang Li, Enmin Feng, and Zhilong Xiu. Parameter identification of nonlinear delayed dynamical system in microbial fermentation based on biological robustness. Numerical Algebra, Control and Optimization, 4(2):103–113, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.103.

Caccetta:2014:MIP

[164] Louis Caccetta and Syarifah Z. Nordin.

Mixed integer programming model for scheduling in unrelated parallel processor system with priority consideration.

Numerical Algebra, Control and Optimization, 4(2):115–132, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.115.

Lee:2014:ECR

[165] Byung-Soo Lee. Existence and convergence results for best proximity points in cone metric spaces. Numerical Algebra, Control and Optimization, 4(2):133–140, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.133.

Kheirfam:2014:WPF

[166] Behrouz Kheirfam. A weighted-path-following method for symmetric cone linear complementarity problems. Numerical Algebra, Control and Optimization, 4(2):141–150, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.141.

Scott:2014:ASD

[167] Jason R. Scott and Stephen Campbell. Auxiliary signal design for failure detection in differential-algebraic equations. Numerical Algebra, Control and Optimization, 4(2):151–179, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.151.

Mahmoud:2014:RCD

[168] Magdi S. Mahmoud and Omar Al-Buraiki. Robust control design of autonomous bicycle kinematics. Numerical Algebra, Control and Optimization, 4(3):181–191, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.181.

Li:2014:CAW

[169] Leong-Kwan Li and Sally Shao. Convergence analysis of the weighted state space search algorithm for recurrent neural networks. Numerical Algebra, Control and Optimization, 4(3):193–207, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.193.

Wu:2014:TSM

[170] Tingting Wu, Yufei Yang, and Huichao Jing. Two-step methods for image zooming using duality strategies. Numerical Algebra, Control and Optimization, 4(3):209–225, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.209.

Lungten:2014:SII

[171] Sangye Lungten, Wil H. A. Schilders, and Joseph M. L. Maubach. Sparse inverse incidence matrices for Schilders’ factorization applied to resistor network modeling. Numerical Algebra, Control and Optimization, 4(3): 227–239, ???? 2014. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.227.

Held:2014:SPA

[172] Harald Held, Gabriela Martinez, and Philipp Emanuel Stelzig. Stochastic programming approach for energy management in electric microgrids. Numerical Algebra, Control and Optimization, 4(3):241–267, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.241.

Aizam:2014:CMT

[173] Nur Aidya Hanum Aizam and Louis Caccetta. Computational models for timetabling problem. Numerical Algebra, Control and Optimization, 4 (3):269–285, ???? 2014. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.269.

Lee:2014:SCI

[174] B. S. Lee and Arif Rafiq. Strong convergence of an implicit iteration process for a finite family of Lipschitz ϕ-uniformly pseudocontractive mappings in Banach spaces. Numerical Algebra, Control and Optimization, 4(4):287–293, ???? 2014. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.287.

Li:2014:NSA

[175] Lili Li and Chunrong Chen. Nonlinear scalarization with applications to Hölder continuity of approximate solutions. Numerical Algebra, Control and Optimization, 4(4):295–307, ???? 2014. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.295.

Yu:2014:TPH

[176] Guolin Yu. Topological properties of Henig globally efficient solutions of set-valued problems. Numerical Algebra, Control and Optimization, 4(4):309–316, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.309.

Qu:2014:NSS

[177] Wei Qu, Siu-Long Lei, and Seak-Weng Vong. A note on the stability of a second order finite difference scheme for space fractional diffusion equations. Numerical Algebra, Control and Optimization, 4(4):317–325, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.317.

Zhang:2014:MPS

[178] Yu Zhang and Tao Chen. Minimax problems for set-valued mappings with set optimization. Numerical Algebra, Control and Optimization, 4(4):327–340, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2014.4.327.

Chan:2014:EIS

[179] Kit Yan Chan, Changjun Yu, Kok Lay Teo, and Sven Nordholm. Essential issues on solving optimal power flow problems using soft-computing. Numerical Algebra, Control and Optimization, 4(4):341–351, ???? 2014. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2014.4.341.

Gao:2015:P

[180] Yan Gao, Zhiqiang Xu, Lei Wang, and Honglei Xu. Preface. Numerical Algebra, Control and Optimization, 5(1):i–ii, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.1i.

Han:2015:DVH

[181] Yanli Han and Yan Gao. Determining the viability for hybrid control systems on a region with piecewise smooth boundary. Numerical Algebra, Control and Optimization, 5(1):1–9, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.1.

Wang:2015:DRD

[182] Li-Min Wang, Jing-Xian Yu, Jia Shi, and Fu-Rong Gao. Delay-range dependent H_{∞} control for uncertain 2d-delayed systems. Numerical Algebra, Control and Optimization, 5 (1):11–23, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.11.

Qi:2015:GCP

[183] Liyan Qi, Xiantao Xiao, and Liwei Zhang. On the global convergence of a parameter-adjusting Levenberg– Marquardt method. Numerical Algebra, Control and Optimization, 5 (1):25–36, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.25.

Li:2015:ACP

[184] Siqi Li and Weiyi Qian. Analysis of complexity of primal-dual interior-point algorithms based on a new kernel function for linear optimization. Numerical Algebra, Control and Optimization, 5(1):37–46, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.37.

Lv:2015:OPT

[185] Wei Lv and Ruirui Sui. Optimality of piecewise thermal conductivity in a snow-ice thermodynamic system. Numerical Algebra, Control and Optimization, 5(1):47–57, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.47.

Zhai:2015:ODS

[186] Jingang Zhai, Guangmao Jiang, and Jianxiong Ye. Optimal dilution strategy for a microbial continuous culture based on the biological robustness. Numerical Algebra, Control and Optimization, 5(1):59–69, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.59.

Luo:2015:SMB

[187] Chengxin Luo. Single machine batch scheduling problem to minimize makespan with controllable setup and jobs processing times. Numerical Algebra, Control and Optimization, 5(1):71–77, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.71.

Liu:2015:PIG

[188] Sanming Liu, Zhijie Wang, and Chongyang Liu. Proximal iterative Gaussian smoothing algorithm for a class of nonsmooth convex minimization problems. Numerical Algebra, Control and Optimization, 5 (1):79–89, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.79.

Bai:2015:P

[189] Yanqin Bai, Duan Li, Hezhi Luo, and Guoqiang Wang. Preface. Numerical Algebra, Control and Optimization, 5(2):i–ii, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.2i.

Wang:2015:PDA

[190] Fengmin Wang, Dachuan Xu, Donglei Du, and Chenchen Wu. Primaldual approximation algorithms for submodular cost set cover problems with linear/submodular penalties. Numerical Algebra, Control and Optimization, 5(2):91–100, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.91.

Wang:2015:CAP

[191] Guoqiang Wang, Zhongchen Wu, Zhongtuan Zheng, and Xinzhong Cai. Complexity analysis of primal-dual interior-point methods for semidefinite optimization based on a parametric kernel function with a trigonometric barrier term. Numerical Algebra, Control and Optimization, 5(2):101–113, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.101.

Li:2015:SQF

[192] Jianling Li, Chunting Lu, and Youfang Zeng. A smooth QP-free algorithm without a penalty function or a filter for mathematical programs with complementarity constraints. Numerical Algebra, Control and Optimization, 5(2):115–126, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.115.

Xiu:2015:ROS

[193] Xianchao Xiu and Lingchen Kong.

Rank-one and sparse matrix decomposition for dynamic MRI. Numerical Algebra, Control and Optimization, 5(2):127–134, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.127.

Msigwa:2015:PBA

[194] Robert Ebihart Msigwa, Yue Lu, Xiantao Xiao, and Liwei Zhang. A perturbation-based approach for continuous network design problem with emissions. Numerical Algebra, Control and Optimization, 5(2):135–149, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.135.

Zhu:2015:SMA

[195] Wenxing Zhu, Yanpo Liu, and Geng Lin. Speeding up a memetic algorithm for the max-bisection problem. Numerical Algebra, Control and Optimization, 5(2):151–168, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.151.

Zhang:2015:WTR

[196] Liang Zhang, Wenyu Sun, Raimundo J. B. de Sampaio, and Jinyun Yuan. A wedge trust region method with self-correcting geometry for derivativefree optimization. Numerical Algebra, Control and Optimization, 5 (2):169–184, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2015.5.169.

Xia:2015:NSR

[197] Yong Xia, Yu-Jun Gong, and Sheng-Nan Han. A new semidefinite relaxation for L_{1}-constrained quadratic optimization and extensions. Numerical Algebra, Control and Optimization, 5(2):185–195, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.185.

Chen:2015:CST

[198] Zhiping Chen and Youpan Han. Continuity and stability of two-stage stochastic programs with quadratic continuous recourse. Numerical Algebra, Control and Optimization, 5 (2):197–209, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.197.

Bai:2015:PDI

[199] Yanqin Bai, Xuerui Gao, and Guoqiang Wang. Primal-dual interiorpoint algorithms for convex quadratic circular cone optimization. Numerical Algebra, Control and Optimization, 5(2):211–231, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.211.

Enkhbat:2015:PPA

[200] R. Enkhbat, N. Tungalag, and A. S.

Strekalovsky. Pseudoconvexity properties of average cost functions. Numerical Algebra, Control and Optimization, 5(3):233–236, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.233.

Zhu:2015:QNT

[201] Honglan Zhu, Qin Ni, and Meilan Zeng.

A quasi-Newton trust region method based on a new fractional model. Numerical Algebra, Control and Optimization, 5(3):237–249, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.237.

Kek:2015:GAO

[202] Sie Long Kek, Mohd Ismail Abd Aziz, and Kok Lay Teo. A gradient algorithm for optimal control problems with model-reality differences. Numerical Algebra, Control and Optimization, 5(3):251–266, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.251.

Mishra:2015:MGM

[203] Debasisha Mishra. Matrix group monotonicity using a dominance notion. Numerical Algebra, Control and Optimization, 5(3):267–274, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.267.

Kek:2015:ORD

[204] Sie Long Kek and Mohd Ismail Abd Aziz. Output regulation for discretetime nonlinear stochastic optimal control problems with model-reality differences. Numerical Algebra, Control and Optimization, 5(3):275–288, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.275.

Tian:2015:SRI

[205] Yongge Tian. A survey on rank and inertia optimization problems of the matrix-valued function A + BXB^{∗}. Numerical Algebra, Control and Optimization, 5(3):289–326, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.289.

Tan:2015:DSS

[206] Xinxin Tan, Shujuan Li, Sisi Liu, Zhiwei Zhao, Lisa Huang, and Jiatai Gang. Dynamic simulation of a SEIQR-v epidemic model based on cellular automata. Numerical Algebra, Control and Optimization, 5(4):327–337, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.327.

Gao:2015:OCM

[207] Jinggui Gao, Xiaoyan Zhao, and Jinggang Zhai. Optimal control of microbial fed-batch culture involving multiple feeds. Numerical Algebra, Control and Optimization, 5(4):339–349, ???? 2015. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.339.

Wang:2015:SSP

[208] Jinzhi Wang and Yuduo Zhang. Solving the seepage problems with free surface by mathematical programming method. Numerical Algebra, Control and Optimization, 5(4):351–357, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.351.

Mao:2015:MIH

[209] Yanan Mao, Caixia Gao, Ruidong Yan, and Aruna Bai. Modeling and identification of hybrid dynamic system in microbial continuous fermentation. Numerical Algebra, Control and Optimization, 5(4):359–368, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.359.

Shen:2015:NSD

[210] Bangyu Shen, Xiaojing Wang, and Chongyang Liu. Nonlinear statedependent impulsive system in fed-batch culture and its optimal control. Numerical Algebra, Control and Optimization, 5(4):369–380, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.369.

Wang:2015:SMM

[211] Yan Wang, Lei Wang, Yanxiang Zhao, Aimin Song, and Yanping Ma. A stochastic model for microbial fermentation process under Gaussian white noise environment. Numerical Algebra, Control and Optimization, 5(4):381–392, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.381.

Zhang:2015:MID

[212] Xu Zhang and Xiang Li. Modeling and identification of dynamical system with genetic regulation in batch fermentation of glycerol. Numerical Algebra, Control and Optimization, 5(4):393–403, ???? 2015. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2015.5.393.

Benner:2016:BBM

[213] Peter Benner, Jens Saak, and M. Monir Uddin. Balancing based model reduction for structured index-2 unstable descriptor systems with application to flow control. Numerical Algebra, Control and Optimization, 6(1):1–20, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016.6.1.

Zhang:2016:OLR

[214] Xuepeng Zhang and Zhibin Liang. Optimal layer reinsurance on the maximization of the adjustment coefficient. Numerical Algebra, Control and Optimization, 6(1):21–34, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016.6.21.

Yu:2016:GPE

[215] Guolin Yu. Global proper efficiency and vector optimization with cone-arcwise connected set-valued maps. Numerical Algebra, Control and Optimization, 6(1):35–44, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016.6.35.

Yan:2016:NCP

[216] Xi-Hong Yan. A new convergence proof of augmented Lagrangian-based method with full Jacobian decomposition for structured variational inequalities. Numerical Algebra, Control and Optimization, 6(1):45–54, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016.6.45.

Liang:2016:DRI

[217] Qiao Liang and Qiang Ye. Deflation by restriction for the inverse-free preconditioned Krylov subspace method. Numerical Algebra, Control and Optimization, 6(1):55–71, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016.6.55.

Hossain:2016:PBM

[218] Mohammad-Sahadet Hossain. Projection-based model reduction for timevarying descriptor systems: New results. Numerical Algebra, Control and Optimization, 6(1):73–90, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016.6.73.

Le:2016:BPN

[219] Thanh Hieu Le and Marc Van Barel. On bounds of the Pythagoras number of the sum of square magnitudes of Laurent polynomials. Numerical Algebra, Control and Optimization, 6(2):91–102, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016001.

Giri:2016:IPN

[220] Chinmay Kumar Giri. Index-proper nonnegative splittings of matrices. Numerical Algebra, Control and Optimization, 6(2):103–113, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016002.

Vieira:2016:DEJ

[221] Manuel V. C. Vieira. Derivatives of eigenvalues and Jordan frames. Numerical Algebra, Control and Optimization, 6(2):115–126, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016003.

Mahmoud:2016:OFO

[222] Magdi S. Mahmoud. Output feedback overlapping control design of interconnected systems with input saturation. Numerical Algebra, Control and Optimization, 6(2):127–151, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016004.

Enkhbat:2016:SMH

[223] Rentsen Enkhbat, M. V. Barkova, and A. S. Strekalovsky. Solving Malfatti’s high dimensional problem by global optimization. Numerical Algebra, Control and Optimization, 6 (2):153–160, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016005.

Sahiner:2016:NSA

[224] Ahmet Sahiner, Gulden Kapusuz, and Nurullah Yilmaz. A new smoothing approach to exact penalty functions for inequality constrained optimization problems. Numerical Algebra, Control and Optimization, 6(2):161–173, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016006.

Hamidoglu:2016:GFT

[225] Ali Hamidoğlu. On general form of the Tanh method and its application to nonlinear partial differential equations. Numerical Algebra, Control and Optimization, 6(2):175–181, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016007.

Kropat:2016:SPD

[226] Erik Kropat, Silja Meyer-Nieberg, and Gerhard-Wilhelm Weber. Singularly perturbed diffusion-advection-reaction processes on extremely large threedimensional curvilinear networks with a periodic microstructure — efficient solution strategies based on homogenization theory. Numerical Algebra, Control and Optimization, 6(2):183–219, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016008.

Sreeram:2016:F

[227] Victor Sreeram. Forword. Numerical Algebra, Control and Optimization, 6(3):i–ii, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.201603i.

Rami:2016:PSH

[228] Mustapha Ait Rami and John Moore.

Partial stabilizability and hidden convexity of indefinite LQ problem. Numerical Algebra, Control and Optimization, 6(3):221–239, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016009.

Helmke:2016:MSR

[229] Uwe Helmke and Michael Schönlein.

Minimum sensitivity realizations of networks of linear systems. Numerical Algebra, Control and Optimization, 6(3):241–262, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016010.

Liang:2016:RFL

[230] A. Liang, C. Wang, W. Liu, and L. Li.

Robust and flexible landmarks detection for uncontrolled frontal faces in the wild. Numerical Algebra, Control and Optimization, 6(3):263–296, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016011.

Li:2016:LCZ

[231] Bin Li, Hai Huyen Dam, and Antonio Cantoni. A low-complexity zero-forcing beamformer design for multiuser MIMO systems via a dual gradient method. Numerical Algebra, Control and Optimization, 6(3):297–304, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016012.

Lee:2016:DCS

[232] Sang-Heon Lee. Development of concurrent structural decentralised discrete event system using bisimulation concept. Numerical Algebra, Control and Optimization, 6(3):305–317, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016013.

Anderson:2016:DGA

[233] Brian D. O. Anderson, Shaoshuai Mou, A. Stephen Morse, and Uwe Helmke. Decentralized gradient algorithm for solution of a linear equation. Numerical Algebra, Control and Optimization, 6(3):319–328, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016014.

Kumar:2016:PFE

[234] Deepak Kumar, Ahmad Jazlan, Victor Sreeram, and Roberto Togneri. Partial fraction expansion based frequency weighted model reduction for discrete-time systems. Numerical Algebra, Control and Optimization, 6 (3):329–337, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016015.

Wang:2016:NCS

[235] Yujing Wang, Changjun Yu, and Kok Lay Teo. A new computational strategy for optimal control problem with a cost on changing control. Numerical Algebra, Control and Optimization, 6(3):339–364, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016016.

Krishnamurthy:2016:IDS

[236] Vikram Krishnamurthy and William Hoiles. Information diffusion in social sensing. Numerical Algebra, Control and Optimization, 6(3):365–411, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016017.

Kramer:2016:PPM

[237] Boris Kramer and John R. Singler.

A POD projection method for largescale algebraic Riccati equations. Numerical Algebra, Control and Optimization, 6(4):413–435, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016018.

Hovda:2016:CFE

[238] Sigve Hovda. Closed-form expression for the inverse of a class of tridiagonal matrices. Numerical Algebra, Control and Optimization, 6(4):437–445, ???? 2016.

CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016019.

Campbell:2016:SHI

[239] Stephen Campbell and Peter Kunkel. Solving higher index DAE optimal control problems. Numerical Algebra, Control and Optimization, 6 (4):447–472, ???? 2016. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016020.

Liu:2016:ASS

[240] Yongchao Liu, Hailin Sun, and Huifu Xu. An approximation scheme for stochastic programs with second order dominance constraints. Numerical Algebra, Control and Optimization, 6(4):473–490, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016021.

Pozdyayev:2016:SAD

[241] Vladimir Pozdyayev. 2d system analysis via dual problems and polynomial matrix inequalities. Numerical Algebra, Control and Optimization, 6(4):491–504, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2016022.

Dang:2016:CAP

[242] Yazheng Dang, Fanwen Meng, and Jie Sun. Convergence analysis of a parallel projection algorithm for solving convex feasibility problems. Numerical Algebra, Control and Optimization, 6(4):505–519, ???? 2016. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2016023.

Glizer:2017:SIH

[243] Valery Y. Glizer and Oleg Kelis. Singular infinite horizon zero-sum linearquadratic differential game: Saddlepoint equilibrium sequence. Numerical Algebra, Control and Optimization, 7(1):1–20, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017001.

Pervin:2017:TEI

[244] Magfura Pervin, Sankar Kumar Roy, and Gerhard Wilhelm Weber. A twoechelon inventory model with stockdependent demand and variable holding cost for deteriorating items. Numerical Algebra, Control and Optimization, 7(1):21–50, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017002.

Kropat:2017:HOC

[245] Erik Kropat. Homogenization of optimal control problems on curvilinear networks with a periodic microstructure — results on S-homogenization and Γ-convergence. Numerical Algebra, Control and Optimization, 7 (1):51–76, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017003.

Eshkuvatov:2017:EAM

[246] Z. K. Eshkuvatov, M. Kammuji, Bachok M. Taib, and N. M. A. Nik Long. Effective approximation method for solving linear Fredholm–Volterra integral equations. Numerical Algebra, Control and Optimization, 7 (1):77–88, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017004.

Ibragimov:2017:SLP

[247] Gafurjan Ibragimov, Askar Rakhmanov, Idham Arif Alias, and Mai Zurwatul Ahlam Mohd Jaffar. The soft landing problem for an infinite system of second order differential equations. Numerical Algebra, Control and Optimization, 7(1):89–94, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017005.

Ibrahim:2017:AOB

[248] Z. B. Ibrahim, N. A. A. Mohd Nasir, K. I. Othman, and N. Zainuddin. Adaptive order of block backward differentiation formulas for stiff ODEs. Numerical Algebra, Control and Optimization, 7(1):95–106, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017006.

Karimi:2017:RBI

[249] Mostafa Karimi, Noor Akma Ibrahim, Mohd Rizam Abu Bakar, and Jayanthi Arasan. Rank-based inference for the accelerated failure time model in the presence of interval censored data. Numerical Algebra, Control and Optimization, 7(1):107–112, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017007.

Ketabchi:2017:CMN

[250] Saeed Ketabchi, Hossein Moosaei, M. Parandegan, and Hamidreza Navidi. Computing minimum norm solution of linear systems of equations by the generalized Newton method. Numerical Algebra, Control and Optimization, 7(2):113–119, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017008.

Veremey:2017:SHO

[251] Evgeny I. Veremey and Vladimir V. Eremeev. SISO h-optimal synthesis with initially specified structure of control law. Numerical Algebra, Control and Optimization, 7(2):121–138, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017009.

Kropat:2017:CNS

[252] Erik Kropat, Silja Meyer-Nieberg, and Gerhard-Wilhelm Weber. Computational networks and systemshomogenization of self-adjoint differential operators in variational form on periodic networks and microarchitectured systems. Numerical Algebra, Control and Optimization, 7 (2):139–169, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017010.

Kheirfam:2017:IFN

[253] Behrouz Kheirfam and Guoqiang Wang. An infeasible full NT-step interior point method for circular optimization. Numerical Algebra, Control and Optimization, 7(2):171–184, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017011.

Shmyrov:2017:OSO

[254] Alexander Shmyrov and Vasily Shmyrov. The optimal stabilization of orbital motion in a neighborhood of collinear libration point. Numerical Algebra, Control and Optimization, 7(2):185–189, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017012.

Srochko:2017:SOC

[255] Vladimir Srochko, Vladimir Antonik, and Elena Aksenyushkina. Sufficient optimality conditions for extremal controls based on functional increment formulas. Numerical Algebra, Control and Optimization, 7(2):191–199, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017013.

Sorokin:2017:FNO

[256] Stepan Sorokin and Maxim Staritsyn. Feedback necessary optimality conditions for a class of terminally constrained state-linear variational problems inspired by impulsive control. Numerical Algebra, Control and Optimization, 7(2):201–210, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017014.

Enkhbat:2017:GOR

[257] Rentsen Enkhbat, Evgeniya A. Finkelstein, Anton S. Anikin, and Alexandr Yu. Gornov. Global optimization reduction of generalized Malfatti’s problem. Numerical Algebra, Control and Optimization, 7(2):211–221, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017015.

Kropat:2017:BGB

[258] Erik Kropat, Silja Meyer-Nieberg, and Gerhard-Wilhelm Weber. Bridging the gap between variational homogenization results and two-scale asymptotic averaging techniques on periodic network structures. Numerical Algebra, Control and Optimization, 7(3):223–250, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017016.

Muradova:2017:ANF

[259] Aliki D. Muradova, Georgios K. Tairidis, and Georgios E. Stavroulakis. Adaptive neuro-fuzzy vibration control of a smart plate. Numerical Algebra, Control and Optimization, 7(3):251–271, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017017.

Li:2017:OOP

[260] Wen Li, Song Wang, and Volker Rehbock. A 2nd-order one-point numerical integration scheme for fractional ordinary differential equations. Numerical Algebra, Control and Optimization, 7(3):273–287, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017018.

Chretien:2017:PCC

[261] Stéphane Chrétien, Sébastien Darses, Christophe Guyeux, and Paul Clarkson. On the pinning controllability of complex networks using perturbation theory of extreme singular values. application to synchronisation in power grids. Numerical Algebra, Control and Optimization, 7(3):289–299, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017019.

Tawhid:2017:SGW

[262] Mohamed A. Tawhid and Ahmed F. Ali. A simplex grey wolf optimizer for solving integer programming and minimax problems. Numerical Algebra, Control and Optimization, 7(3):301–323, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2017020.

Yang:2017:AOB

[263] Wanli Yang, Jie Sun, and Su Zhang.

Analysis of optimal boundary control for a three-dimensional reaction-diffusion system. Numerical Algebra, Control and Optimization, 7(3):325–344, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017021.

Zhang:2017:TNC

[264] Yibo Zhang, Jinfeng Gao, Jia Ren, and Huijiao Wang. A type of new consensus protocol for two-dimension multi-agent systems. Numerical Algebra, Control and Optimization, 7 (3):345–357, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017022.

Ahmadi:2017:MSP

[265] Ardeshir Ahmadi and Hamed Davari-Ardakani. A multistage stochastic programming framework for cardinality constrained portfolio optimization. Numerical Algebra, Control and Optimization, 7(3):359–377, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017023.

Moslemi:2017:PEF

[266] Shiva Moslemi and Abolfazl Mirzazadeh. Performance evaluation of fourstage blood supply chain with feedback variables using NDEA cross-efficiency and entropy measures under IER uncertainty. Numerical Algebra, Control and Optimization, 7(4):379–401, ???? 2017.

CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017024.

Zahedi-Seresht:2017:NMC

[267] Mazyar Zahedi-Seresht, Gholam-Reza Jahanshahloo, Josef Jablonsky, and Sedighe Asghariniya. A new Monte Carlo based procedure for complete ranking efficient units in DEA models. Numerical Algebra, Control and Optimization, 7(4):403–416, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017025.

Tirkolaee:2017:RMT

[268] Erfan Babaee Tirkolaee, Alireza Goli, Mani Bakhsi, and Iraj Mahdavi. A robust multi-trip vehicle routing problem of perishable products with intermediate depots and time windows. Numerical Algebra, Control and Optimization, 7(4):417–433, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017026.

Ahmadzadeh:2017:IMI

[269] Farzaneh Ahmadzadeh, Kathrina Jederström, Maria Plahn, Anna Olsson, and Isabell Foyer. An investigation of the most important factors for sustainable product development using evidential reasoning. Numerical Algebra, Control and Optimization, 7(4):435–455, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017027.

Xiao:2017:SNI

[270] Hongguang Xiao, Wen Tan, Dehua Xiang, Lifu Chen, and Ning Li. A study of numerical integration based on Legendre polynomial and RLS algorithm. Numerical Algebra, Control and Optimization, 7(4):457–464, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017028.

Ardakan:2017:HMH

[271] Mostafa Abouei Ardakan, A. Kourank Beheshti, S. Hamid Mirmohammadi, and Hamed Davari Ardakani. A hybrid meta-heuristic algorithm to minimize the number of tardy jobs in a dynamic two-machine flow shop problem. Numerical Algebra, Control and Optimization, 7(4):465–480, ???? 2017. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017029.

Tyatyushkin:2017:NMS

[272] Alexander Tyatyushkin and Tatiana Zarodnyuk. Numerical method for solving optimal control problems with phase constraints. Numerical Algebra, Control and Optimization, 7 (4):481–492, ???? 2017. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2017030.

Filar:2018:LGR

[273] Jerzy A. Filar, Michael Haythorpe, and Richard Taylor. Linearly-growing reductions of Karp’s 21 NP-complete problems. Numerical Algebra, Control and Optimization, 8(1):1–16, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018001.

Csomos:2018:FSM

[274] Petra Csomós and Hermann Mena. Fourier-splitting method for solving hyperbolic LQR problems. Numerical Algebra, Control and Optimization, 8(1):17–46, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018002.

Gergel:2018:GNS

[275] Victor Gergel, Konstantin Barkalov, and Alexander Sysoyev. Globalizer: A novel supercomputer software system for solving time-consuming global optimization problems. Numerical Algebra, Control and Optimization, 8(1):47–62, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018003.

Mostolizadeh:2018:MMC

[276] Reihaneh Mostolizadeh, Zahra Afsharnezhad, and Anna Marciniak-Czochra. Mathematical model of Chimeric Anti-gene Receptor (CAR) T cell therapy with presence of cytokine. Numerical Algebra, Control and Optimization, 8(1):63–80, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018004.

Adly:2018:USD

[277] Samir Adly and Ba Khiet Le. Unbounded state-dependent sweeping processes with perturbations in uniformly convex and q-uniformly smooth Banach spaces. Numerical Algebra, Control and Optimization, 8(1):81–95, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018005.

Liu:2018:FLP

[278] Yanqing Liu, Jiyuan Tao, Huan Zhang, Xianchao Xiu, and Lingchen Kong. Fused LASSO penalized least absolute deviation estimator for high dimensional linear regression. Numerical Algebra, Control and Optimization, 8(1):97–117, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018006.

Benner:2018:NCL

[279] Peter Benner, Ryan Lowe, and Matthias Voigt. L_{∞}-norm computation for largescale descriptor systems using structured iterative eigensolvers. Numerical Algebra, Control and Optimization, 8(1):119–133, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018007.

Kropat:2018:FTE

[280] Erik Kropat and Gerhard Wilhelm Weber. Fuzzy target-environment networks and fuzzy-regression approaches. Numerical Algebra, Control and Optimization, 8(2):135–155, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018008.

Ledari:2018:TER

[281] Ashkan Mohsenzadeh Ledari, Alireza Arshadi Khamseh, and Mohammad Mohammadi. A three echelon revenue oriented green supply chain network design. Numerical Algebra, Control and Optimization, 8(2):157–168, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018009.

Pervin:2018:IIM

[282] Magfura Pervin, Sankar Kumar Roy, and Gerhard Wilhelm Weber. An integrated inventory model with variable holding cost under two levels of trade-credit policy. Numerical Algebra, Control and Optimization, 8(2):169–191, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018010.

Arguchintsev:2018:OCP

[283] Alexander Arguchintsev and Vasilisa Poplevko. An optimal control problem by parabolic equation with boundary smooth control and an integral constraint. Numerical Algebra, Control and Optimization, 8(2):193–202, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018011.

Gu:2018:NBE

[284] Yining Gu and Wei Wu. New bounds for eigenvalues of strictly diagonally dominant tensors. Numerical Algebra, Control and Optimization, 8(2):203–210, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018012.

Glizer:2018:API

[285] Valery Y. Glizer and Oleg Kelis. Asymptotic properties of an infinite horizon partial cheap control problem for linear systems with known disturbances. Numerical Algebra, Control and Optimization, 8(2):211–235, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018013.

Duff:2018:ESC

[286] Iain Duff, Jonathan Hogg, and Florent Lopez. Experiments with sparse Cholesky using a sequential taskflow implementation. Numerical Algebra, Control and Optimization, 8 (2):237–260, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018014.

Kheirfam:2018:EAS

[287] Behrouz Kheirfam and Morteza Moslemi. On the extension of an arcsearch interior-point algorithm for semidefinite optimization. Numerical Algebra, Control and Optimization, 8(2):261–275, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018015.

Chen:2018:P

[288] Chuei Yee Chen and Lai Soon Lee. Preface. Numerical Algebra, Control and Optimization, 8(3):i, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.201803i.

Rakhmatillo:2018:CRA

[289] Aloev Rakhmatillo, Khudoyberganov Mirzoali, and Blokhin Alexander. Construction and research of adequate computational models for quasilinear hyperbolic systems. Numerical Algebra, Control and Optimization, 8(3):277–289, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018017.

Ng:2018:PPO

[290] Teck Wee Ng and Siti Nur Iqmal Ibrahim. Pricing down-and-out power options with exponentially curved barrier. Numerical Algebra, Control and Optimization, 8(3):291–297, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2018018.

Ahmed:2018:CTS

[291] Shohel Ahmed, Abdul Alim, and Sumaiya Rahman. A controlled treatment strategy applied to HIV immunology model. Numerical Algebra, Control and Optimization, 8(3):

299–314, ????  2018.  CODEN

???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018019.

Lee:2018:TPA

[292] M. S. Lee, B. S. Goh, H. G. Harno, and K. H. Lim. On a two-phase approximate greatest descent method for nonlinear optimization with equality constraints. Numerical Algebra, Control and Optimization, 8(3):315–326, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018020.

Lim:2018:AGD

[293] King Hann Lim, Hong Hui Tan, and Hendra G. Harno. Approximate greatest descent in neural network optimization. Numerical Algebra, Control and Optimization, 8(3):327–336, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018021.

Eshkuvatov:2018:HPM

[294] Zainidin Eshkuvatov. Homotopy perturbation method and Chebyshev polynomials for solving a class of singular and hypersingular integral equations. Numerical Algebra, Control and Optimization, 8(3):337–350, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018022.

Buba:2018:DEI

[295] Ahmed Tarajo Buba and Lai Soon Lee. Differential evolution with improved sub-route reversal repair mechanism for multiobjective urban transit routing problem. Numerical Algebra, Control and Optimization, 8 (3):351–376, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018023.

Sim:2018:MSS

[296] Hong Seng Sim, Wah June Leong, Chuei Yee Chen, and Siti Nur Iqmal Ibrahim. Multi-step spectral gradient methods with modified weak secant relation for large scale unconstrained optimization. Numerical Algebra, Control and Optimization, 8(3):377–387, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018024.

Li:2018:PPC

[297] Xing Li, Chungen Shen, and Lei-Hong Zhang. A projected preconditioned conjugate gradient method for the linear response eigenvalue problem. Numerical Algebra, Control and Optimization, 8(4):389–412, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018025.

Arasu:2018:OPO

[298] K. T. Arasu and Manil T. Mohan. Optimization problems with orthogonal matrix constraints. Numerical Algebra, Control and Optimization, 8(4):413–440, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018026.

Khatibzadeh:2018:CPP

[299] Hadi Khatibzadeh, Vahid Mohebbi, and Mohammad Hossein Alizadeh. On the cyclic pseudomonotonicity and the proximal point algorithm. Numerical Algebra, Control and Optimization, 8(4):441–449, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018027.

Liu:2018:QSA

[300] Yongchao Liu. Quantitative stability analysis of stochastic mathematical programs with vertical complementarity constraints. Numerical Algebra, Control and Optimization, 8(4):451–460, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018028.

Dolatabadi:2018:WVO

[301] Soheil Dolatabadi. Weighted vertices optimizer (WVO): A novel metaheuristic optimization algorithm. Numerical Algebra, Control and Optimization, 8(4):461–479, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018029.

Yilmaz:2018:SAM

[302] Bilgi Yilmaz and A. Sevtap Selcuk-Kestel. A stochastic approach to model housing markets: The US housing market case. Numerical Algebra, Control and Optimization, 8 (4):481–492, ???? 2018. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018030.

Ma:2018:FRP

[303] Haifeng Ma and Xiaoshuang Gao. Further results on the perturbation estimations for the Drazin inverse. Numerical Algebra, Control and Optimization, 8(4):493–503, ???? 2018. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2018031.

Mohammad:2019:BSM

[304] Hassan Mohammad, Mohammed Yusuf Waziri, and Sandra Augusta Santos. A brief survey of methods for solving nonlinear least-squares problems. Numerical Algebra, Control and Optimization, 9(1):1–13, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019001.

Guerarra:2019:PND

[305] Sihem Guerarra. Positive and negative definite submatrices in an Hermitian least rank solution of the matrix equation AXA^{∗} = B. Numerical Algebra, Control and Optimization, 9(1):15–22, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019002.

Pulch:2019:SPG

[306] Roland Pulch. Stability preservation in Galerkin-type projection-based model order reduction. Numerical Algebra, Control and Optimization, 9(1):23–44, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019003.

Qian:2019:LSR

[307] Rui Qian, Rong Hu, and Ya-Ping Fang.

Local smooth representation of solution sets in parametric linear fractional programming problems. Numerical Algebra, Control and Optimization, 9(1):45–52, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019004.

Pavlikova:2019:CUL

[308] Soña Pavlíková and Daniel Ševčovič. On construction of upper and lower bounds for the HOMO-LUMO spectral gap. Numerical Algebra, Control and Optimization, 9(1):53–69, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019005.

Zhai:2019:FOI

[309] Wenjuan Zhai and Bingzhen Chen.

A fourth order implicit symmetric and symplectic exponentially fitted Runge–Kutta–Nyström method for solving oscillatory problems. Numerical Algebra, Control and Optimization, 9(1):71–84, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019006.

Zhong:2019:SLC

[310] Hongxiu Zhong, Guoliang Chen, and Xueping Guo. Semi-local convergence of the Newton-HSS method under the center Lipschitz condition. Numerical Algebra, Control and Optimization, 9(1):85–99, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019007.

Kheirabadi:2019:SOC

[311] Akram Kheirabadi, Asadollah Mahmoudzadeh Vaziri, and Sohrab Effati. Solving optimal control problem using Hermite wavelet. Numerical Algebra, Control and Optimization, 9(1):101–112, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019008.

Liu:2019:IMO

[312] Qiong Liu, Ahmad Reza Rezaei, Kuan Yew Wong, and Mohammad Mahdi Azami. Integrated modeling and optimization of material flow and financial flow of supply chain network considering financial ratios. Numerical Algebra, Control and Optimization, 9(2):113–132, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019009.

Jayswala:2019:SOM

[313] Anurag Jayswala, Tadeusz Antczakb, and Shalini Jha. Second order modified objective function method for twice differentiable vector optimization problems over cone constraints. Numerical Algebra, Control and Optimization, 9(2):133–145, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019010.

Asadi:2019:MTP

[314] Soodabeh Asadi and Hossein Mansouri. A Mehrotra type predictorcorrector interior-point algorithm for linear programming. Numerical Algebra, Control and Optimization, 9 (2):147–156, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019011.

Abaffy:2019:NRC

[315] József Abaffy. A new reprojection of the conjugate directions. Numerical Algebra, Control and Optimization, 9(2):157–171, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019012.

Hossain:2019:IMS

[316] M. Sumon Hossain and M. Monir Uddin. Iterative methods for solving large sparse Lyapunov equations and application to model reduction of index 1 differential-algebraic-equations. Numerical Algebra, Control and Optimization, 9(2):173–186, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019013.

Goli:2019:ARO

[317] Alireza Goli, Hasan Khademi Zare, Reza Tavakkoli-Moghaddam, and Ahmad Sadeghieh. Application of robust optimization for a product portfolio problem using an invasive weed optimization algorithm. Numerical Algebra, Control and Optimization, 9(2):187–209, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019014.

Dong:2019:HMM

[318] Zhengshan Dong, Jianli Chen, and Wenxing Zhu. Homotopy method for matrix rank minimization based on the matrix hard thresholding method. Numerical Algebra, Control and Optimization, 9(2):211–224, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019015.

Chyba:2019:IMF

[319] Monique Chyba and Geoff Patterson. Indirect methods for fuel-minimal rendezvous with a large population of temporarily captured orbiters. Numerical Algebra, Control and Optimization, 9(2):225–256, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019016.

Sreeram:2019:P

[320] Victor Sreeram. Preface. Numerical Algebra, Control and Optimization, 9(3):i, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.201903i.

Chen:2019:BRF

[321] Liangming Chen, Ming Cao, and Chuanjiang Li. Bearing rigidity and formation stabilization for multiple rigid bodies in SE(3). Numerical Algebra, Control and Optimization, 9(3):257–267, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019017.

Wang:2019:RCC

[322] Xuan Wang, Shaoshuai Mou, and Shreyas Sundaram. A resilient convex combination for consensus-based distributed algorithms. Numerical Algebra, Control and Optimization, 9 (3):269–281, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019018.

Argha:2019:OSO

[323] Ahmadreza Argha, Steven W. Su, Lin Ye, and Branko G. Celler. Optimal sparse output feedback for networked systems with parametric uncertainties. Numerical Algebra, Control and Optimization, 9(3):283–295, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019019.

Sugie:2019:IHM

[324] Zhiyong Sun and Toshiharu Sugie. Identification of Hessian matrix in distributed gradient-based multi-agent coordination control systems. Numerical Algebra, Control and Optimization, 9(3):297–318, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019020.

Jazlan:2019:FIM

[325] Ahmad Jazlan, Umair Zulfiqar, Victor Sreeram, Deepak Kumar, Roberto Togneri, and Hasan Firdaus Mohd Zaki. Frequency interval model reduction of complex fir digital filters. Numerical Algebra, Control and Optimization, 9(3):319–326, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019021.

Nguyen:2019:DOA

[326] Chuong Van Nguyen, Phuong Huu Hoang, and Hyo-Sung Ahn. Distributed optimization algorithms for game of power generation in smart grid. Numerical Algebra, Control and Optimization, 9(3):327–348, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019022.

Adibzadeh:2019:COC

[327] Amir Adibzadeh, Mohsen Zamani, Amir A. Suratgar, and Mohammad B. Menhaj. Constrained optimal consensus in dynamical networks. Numerical Algebra, Control and Optimization, 9(3):349–360, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019023.

Achanta:2019:OSP

[328] Hema K. Achanta, Soura Dasgupta, Raghuraman Mudumbai, Weiyu Xu, and Zhi Ding. Optimum sensor placement for localization of a hazardous source under log normal shadowing. Numerical Algebra, Control and Optimization, 9(3):361–382, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019024.

Deistler:2019:SAS

[329] Manfred Deistler. Singular ARMA systems: a structure theory. Numerical Algebra, Control and Optimization, 9(3):383–391, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019025.

Dutta:2019:NSA

[330] Praveen Kumar Gupta and Ajoy Dutta.

Numerical solution with analysis of HIV/AIDS dynamics model with effect of fusion and cure rate. Numerical Algebra, Control and Optimization, 9(4):393–399, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019038.

Ahangari:2019:CDS

[331] Fatemeh Ahangari. Conformal deformations of a specific class of Lorentzian manifolds with non-irreducible holonomy representation. Numerical Algebra, Control and Optimization, 9 (4):401–412, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019039.

Abidin:2019:OBM

[332] Nurul Hafizah Zainal Abidin, Nor Fadzillah Mohd Mokhtar, and Zanariah Abdul Majid. Onset of Benard– Marangoni instabilities in a double diffusive binary fluid layer with temperaturedependent viscosity. Numerical Algebra, Control and Optimization, 9 (4):413–421, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019040.

Najib:2019:SAS

[333] Najwa Najib, Norfifah Bachok, Norihan Md Arifin, and Fadzilah Md Ali. Stability analysis of stagnation point flow in nanofluid over stretching/shrinking sheet with slip effect using Buongiorno’s model. Numerical Algebra, Control and Optimization, 9(4):423–431, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019041.

Rabiei:2019:NSV

[334] Faranak Rabiei, Fatin Abd Hamid, Zanariah Abd Majid, and Fudziah Ismail. Numerical solutions of Volterra integro-differential equations using General Linear Method. Numerical Algebra, Control and Optimization, 9 (4):433–444, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019042.

Salahuddin:2019:SGM

[335] Salahuddin. System of generalized mixed nonlinear ordered variational inclusions. Numerical Algebra, Control and Optimization, 9 (4):445–460, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019026.

Feng:2019:STD

[336] Yang Wang and Yi fu Feng. θ scheme with two dimensional wavelet-like incremental unknowns for a class of porous medium diffusion-type equations. Numerical Algebra, Control and Optimization, 9(4):461–481, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019027.

Salkuyeh:2019:PSI

[337] Tahereh Salimi Siahkolaei and Davod Khojasteh Salkuyeh. A preconditioned SSOR iteration method for solving complex symmetric system of linear equations. Numerical Algebra, Control and Optimization, 9(4):483–492, ???? 2019. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019033.

Alipour:2019:HPA

[338] M. Alipour, M. A. Vali, and A. H.

Borzabadi. A hybrid parametrization approach for a class of nonlinear optimal control problems. Numerical Algebra, Control and Optimization, 9(4):493–506, ???? 2019. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019037.

Yuan:2020:AIS

[339] Ye Yuan, Yan Ren, Xiaodong Liu, and Jing Wang. Approach to image segmentation based on interval neutrosophic set. Numerical Algebra, Control and Optimization, 10(1):1–11, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019028.

Ketabchi:2020:NCS

[340] Fakhrodin Hashemi and Saeed Ketabchi. Numerical comparisons of smoothing functions for optimal correction of an infeasible system of absolute value equations. Numerical Algebra, Control and Optimization, 10(1):13–21, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019029.

Wang:2020:NSO

[341] Song Wang. Numerical solution of an obstacle problem with interval coefficients. Numerical Algebra, Control and Optimization, 10(1):23–38, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019030.

Betts:2020:IGS

[342] John T. Betts, Stephen L. Campbell, and Claire Digirolamo. Initial guess sensitivity in computational optimal control problems. Numerical Algebra, Control and Optimization, 10(1):39–41, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019031.

Roy:2020:IIP

[343] Sankar Kumar Roy, Magfura Pervin, and Gerhard Wilhelm Weber. Imperfection with inspection policy and variable demand under trade-credit: a deteriorating inventory model. Numerical Algebra, Control and Optimization, 10(1):45–74, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019032.

Chadli:2020:EIA

[344] Ouayl Chadli, Gayatri Pany, and Ram N. Mohapatra. Existence and iterative approximation method for solving mixed equilibrium problem under generalized monotonicity in Banach spaces. Numerical Algebra, Control and Optimization, 10(1):75–92, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019034.

Treanta:2020:CES

[345] Savin Treanţă. Characterization of efficient solutions for a class of PDEconstrained vector control problems.

Numerical Algebra, Control and Optimization, 10(1):93–106, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019035.

Wei:2020:UVQ

[346] Hong-Zhi Wei, Xin Zuo, and Chun-Rong Chen. Unified vector quasiequilibrium problems via improvement sets and nonlinear scalarization with stability analysis. Numerical Algebra, Control and Optimization, 10(1):107–125, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019036.

Sadeghi:2020:RAT

[347] Jafar Sadeghi, Mojtaba Ghiyasi, and Akram Dehnokhalaji. Resource allocation and target setting based on virtual profit improvement. Numerical Algebra, Control and Optimization, 10(2):127–142, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019043.

Kheirabadi:2020:LOC

[348] Akram Kheirabadi, Asadollah Mahmoudzadeh Vaziri, and Sohrab Effati. Linear optimal control of time delay systems via Hermite wavelet. Numerical Algebra, Control and Optimization, 10(2):143–156, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019044.

Sun:2020:QIM

[349] Yu-Feng Sun, Zheng Zeng, and Jie Song. Quasilinear iterative method for the boundary value problem of nonlinear fractional differential equation. Numerical Algebra, Control and Optimization, 10(2):157–164, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019045.

Salahi:2020:QOT

[350] Saeid Ansary Karbasy and Maziar Salahi. Quadratic optimization with two ball constraints. Numerical Algebra, Control and Optimization, 10 (2):165–175, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019046.

Tsachouridis:2020:FAS

[351] Vassilios A. Tsachouridis, Georgios Giantamidis, Stylianos Basagiannis, and Kostas Kouramas. Formal analysis of the Schulz matrix inversion algorithm: a paradigm towards computer aided verification of general matrix flow solvers. Numerical Algebra, Control and Optimization, 10(2):177–206, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2019047.

Allali:2020:OCH

[352] Jaouad Danane and Karam Allali. Optimal control of an HIV model with CTL cells and latently infected cells. Numerical Algebra, Control and Optimization, 10(2):207–225, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019048.

Hassan:2020:NTQ

[353] Basim A. Hassan. A new type of quasi-Newton updating formulas based on the new quasi-Newton equation. Numerical Algebra, Control and Optimization, 10(2):227–235, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019049.

Vinodkumar:2020:SIT

[354] K. Aruna Sakthi and A. Vinodkumar. Stabilization on input time-varying delay for linear switched systems with truncated predictor control. Numerical Algebra, Control and Optimization, 10(2):237–247, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019050.

Shipilevsky:2020:CQO

[355] Yuly Shipilevsky. Complex and quaternionic optimization. Numerical Algebra, Control and Optimization, 10 (3):249–255, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2019051.

Asghariniya:2020:RAC

[356] Sedighe Asghariniya, Hamed Zhiani Rezai, and Saeid Mehrabian. Resource allocation: a common set of weights model. Numerical Algebra, Control and Optimization, 10(3):257–273, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020001.

Ghanem:2020:NSB

[357] Radouen Ghanem and Billel Zireg. Numerical solution of bilateral obstacle optimal control problem, where the controls and the obstacles coincide. Numerical Algebra, Control and Optimization, 10(3):275–300, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020002.

Zare:2020:FQO

[358] Arezu Zare, Mohammad Keyanpour, and Maziar Salahi. On fractional quadratic optimization problem with two quadratic constraints. Numerical Algebra, Control and Optimization, 10(3):301–315, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020003.

Yilmaz:2020:NST

[359] Nurullah Yilmaz and Ahmet Sahiner.

On a new smoothing technique for nonsmooth, non-convex optimization. Numerical Algebra, Control and Optimization, 10(3):317–330, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020004.

Xiong:2020:OBC

[360] Junlin Xiong and Wenjie Liu. H_{∞} observer-based control for large-scale systems with sparse observer communication network. Numerical Algebra, Control and Optimization, 10 (3):331–343, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020005.

Li:2020:SMC

[361] Yuan Li, Ruxia Zhang, Yi Zhang, and Bo Yang. Sliding mode control for uncertain T-S fuzzy systems with input and state delays. Numerical Algebra, Control and Optimization, 10(3):345–354, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020006.

Liu:2020:IAT

[362] Zhi Liu and Tie Zhang. An improved ARMA(1, 1) type fuzzy time series applied in predicting disordering. Numerical Algebra, Control and Optimization, 10(3):355–366, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020007.

Liu:2020:FEO

[363] Lei Liu, Shaoying Lu, Cunwu Han, Chao Li, and Zejin Feng. Fault estimation and optimization for uncertain disturbed singularly perturbed systems with timedelay. Numerical Algebra, Control and Optimization, 10(3):367–379, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020008.

Qiao:2020:PCC

[364] Jinglai Qiao, Li Yang, and Jiawei Yao.

Passive control for a class of nonlinear systems by using the technique of adding a power integrator. Numerical Algebra, Control and Optimization, 10(3):381–389, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020009.

Meng:2020:BAS

[365] Xin-You Meng, Yu-Qian Wu, and Jie Li. Bifurcation analysis of a Singular Nutrient–plankton–fish model with taxation, protected zone and multiple delays. Numerical Algebra, Control and Optimization, 10(3):391–423, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020010.

Dai:2020:P

[366] Yu-Hong Dai, Yiju Wang, and Naihua Xiu. Preface. Numerical Algebra, Control and Optimization, 10 (4):i–ii, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020041.

Tong:2020:NSP

[367] Wanbin Tong, Hongjin He, Chen Ling, and Liqun Qi. A nonmonotone spectral projected gradient method for tensor eigenvalue complementarity problems. Numerical Algebra, Control and Optimization, 10(4):425–437, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020042.

Xu:2020:BCT

[368] Zhuoyi Xu, Yong Xia, and Deren Han. On box-constrained total least squares problem. Numerical Algebra, Control and Optimization, 10 (4):439–449, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020043.

Chang:2020:IAG

[369] Xiao-Wen Chang and David Titley-Peloquin. An improved algorithm for generalized least squares estimation. Numerical Algebra, Control and Optimization, 10(4):451–461, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020044.

Hu:2020:CRD

[370] Leyu Hu and Xingju Cai. Convergence of a randomized Douglas– Rachford method for linear system. Numerical Algebra, Control and Optimization, 10(4):463–474, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020045.

Chen:2020:TRA

[371] Yannan Chen and Jingya Chang. A trust region algorithm for computing extreme eigenvalues of tensors. Numerical Algebra, Control and Optimization, 10(4):475–485, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020046.

Gu:2020:ADM

[372] Yan Gu and Nobuo Yamashita. Alternating direction method of multipliers with variable metric indefinite proximal terms for convex optimization. Numerical Algebra, Control and Optimization, 10(4):487–510, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020047.

Li:2020:PRP

[373] Chengjin Li. Parameter-related projection-based iterative algorithm for a kind of generalized positive semidefinite least squares problem. Numerical Algebra, Control and Optimization, 10(4):511–520, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020048.

Zhou:2020:TSS

[374] Bin Zhou and Hailin Sun. Two-stage stochastic variational inequalities for Cournot–Nash equilibrium with riskaverse players under uncertainty. Numerical Algebra, Control and Optimization, 10(4):521–535, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020049.

Xi:2020:SDF

[375] Min Xi, Wenyu Sun, and Jun Chen. Survey of derivative-free optimization. Numerical Algebra, Control and Optimization, 10(4):537–555, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020050.

Yang:2020:PGS

[376] Xin Yang, Nan Wang, and Lingling Xu. A parallel Gauss–Seidel method for convex problems with separable structure. Numerical Algebra, Control and Optimization, 10(4):557–570, ???? 2020. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020051.

Zhang:2020:RPF

[377] Li Zhang, Xiaofeng Zhou, and Min Chen. The research on the properties of Fourier matrix and bent function. Numerical Algebra, Control and Optimization, 10(4):571–578, ???? 2020. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020052.

Zhang:2021:FTC

[378] Xuefeng Zhang and Yingbo Zhang. Fault-tolerant control against actuator failures for uncertain singular fractional order systems. Numerical Algebra, Control and Optimization, 11(1):1–12, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020011.

Yu:2021:DCP

[379] Peizhao Yu, Guoshan Zhang, and Yi Zhang. Decoupling of cubic polynomial matrix systems. Numerical Algebra, Control and Optimization, 11(1):13–26, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020012.

Bastani:2021:GIM

[380] Mehdi Bastani and Davod Khojasteh Salkuyeh. On the GSOR iteration method for image restoration. Numerical Algebra, Control and Optimization, 11(1):27–43, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020013.

Lee:2021:BBC

[381] M. S. Lee, H. G. Harno, B. S. Goh, and K. H. Lim. On the bang-bang control approach via a component-wise line search strategy for unconstrained optimization. Numerical Algebra, Control and Optimization, 11(1):45–61, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020014.

Guerarra:2021:MMR

[382] Sihem Guerarra. Maximum and minimum ranks and inertias of the Hermitian parts of the least rank solution of the matrix equation AXB = C. Numerical Algebra, Control and Optimization, 11(1):75–86, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020016.

Kahya:2021:IWO

[383] Mohammed Abdulrazaq Kahya, Suhaib Abduljabbar Altamir, and Zakariya Yahya Algamal. Improving whale optimization algorithm for feature selection with a time-varying transfer function. Numerical Algebra, Control and Optimization, 11(1):87–98, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020017.

Upadhyaya:2021:DMA

[384] Parikshit Upadhyaya, Elias Jarlebring, and Emanuel H. Rubensson. A density matrix approach to the convergence of the self-consistent field iteration. Numerical Algebra, Control and Optimization, 11(1):99–115, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020018.

Niu:2021:PCM

[385] Hong Niu, Zhijiang Feng, Qijin Xiao, and Yajun Zhang. A PID control method based on optimal control strategy. Numerical Algebra, Control and Optimization, 11(1):117–126, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020019.

Lv:2021:DCU

[386] Hui Lv and Xing’an Wang. Dissipative control for uncertain singular Markovian jump systems via hybrid impulsive control. Numerical Algebra, Control and Optimization, 11(1):127–142, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020020.

Putkaradze:2021:NSR

[387] Vakhtang Putkaradze and Stuart Rogers. Numerical simulations of a rolling ball robot actuated by internal point masses. Numerical Algebra, Control and Optimization, 11 (2):143–207, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020021.

Rentsen:2021:GNE

[388] Enkhbat Rentsen and Battur Gompil.

Generalized Nash equilibrium problem based on Malfatti’s problem. Numerical Algebra, Control and Optimization, 11(2):209–220, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020022.

Lotfi:2021:ROM

[389] Reza Lotfi, Yahia Zare Mehrjerdi, Mir Saman Pishvaee, Ahmad Sadeghieh, and Gerhard-Wilhelm Weber. A robust optimization model for sustainable and resilient closed-loop supply chain network design considering conditional value at risk. Numerical Algebra, Control and Optimization, 11(2):221–253, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020023.

Jahan:2021:DAR

[390] Sohana Jahan. Discriminant analysis of regularized multidimensional scaling. Numerical Algebra, Control and Optimization, 11(2):255–267, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020024.

Uzunca:2021:ADG

[391] Murat Uzunca and Ayşe Sarıaydın-Filibelioğlu. Adaptive discontinuous Galerkin finite elements for advective Allen–Cahn equation. Numerical Algebra, Control and Optimization, 11(2):269–281, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020025.

Betts:2021:ESO

[392] John T. Betts, Stephen Campbell, and Claire Digirolamo. Examination of solving optimal control problems with delays using GPOPS-II. Numerical Algebra, Control and Optimization, 11(2):283–305, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020026.

Glizer:2021:NCE

[393] Valery Y. Glizer. Novel conditions of Euclidean space controllability for singularly perturbed systems with input delay. Numerical Algebra, Control and Optimization, 11(2):307–320, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020027.

Abdolhosseinzadeh:2021:DET

[394] Mohsen Abdolhosseinzadeh and Mir Mohammad Alipour. Design of experiment for tuning parameters of an ant colony optimization method for the constrained shortest Hamiltonian path problem in the grid networks. Numerical Algebra, Control and Optimization, 11(2):321–332, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020028.

Ghelichi:2021:NFA

[395] Mostafa Ghelichi, A. M. Goltabar, H. R.

Tavakoli, and A. Karamodin. Neurofuzzy active control optimized by tug of war optimization method for seismically excited benchmark highway bridge. Numerical Algebra, Control and Optimization, 11(3):333–351, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020029.

He:2021:IAD

[396] Jie-Wen He, Chi-Chon Lei, Chen-Yang Shi, and Seak-Weng Vong. An inexact alternating direction method of multipliers for a kind of nonlinear complementarity problems. Numerical Algebra, Control and Optimization, 11(3):353–362, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020030.

Danane:2021:OCV

[397] Jaouad Danane. Optimal control of viral infection model with saturated infection rate. Numerical Algebra, Control and Optimization, 11(3):363–375, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020031.

Delladji:2021:BCP

[398] Sarra Delladji, Mohammed Belloufi, and Badreddine Sellami. Behavior of the combination of PRP and HZ methods for unconstrained optimization. Numerical Algebra, Control and Optimization, 11(3):377–389, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020032.

Panja:2021:DSS

[399] Prabir Panja, Soovoojeet Jana, and Shyamal kumar Mondal. Dynamics of a stage structure prey–predator model with ratio-dependent functional response and anti-predator behavior of adult prey. Numerical Algebra, Control and Optimization, 11 (3):391–405, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020033.

Breiten:2021:SDR

[400] Tobias Breiten, Sergey Dolgov, and Martin Stoll. Solving differential Riccati equations: a nonlinear space–time method using tensor trains. Numerical Algebra, Control and Optimization, 11(3):407–429, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020034.

Noor:2021:PHO

[401] Muhammad Aslam Noor and Khalida Inayat Noor. Properties of higher order preinvex functions. Numerical Algebra, Control and Optimization, 11(3):431–441, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020035.

Rentsen:2021:AST

[402] Enkhbat Rentsen, N. Tungalag, J. Enkhbayar, O. Battogtokh, and L. Enkhtuvshin. Application of survival theory in mining industry. Numerical Algebra, Control and Optimization, 11(3):443–448, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020036.

Shahsavari:2021:PMS

[403] Samira Shahsavari and Saeed Ketabchi.

The proximal methods for solving absolute value equation. Numerical Algebra, Control and Optimization, 11(3):449–460, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020037.

Ravikumar:2021:ACN

[404] K. Ravikumar, Manil T. Mohan, and A. Anguraj. Approximate controllability of a non-autonomous evolution equation in Banach spaces. Numerical Algebra, Control and Optimization, 11(3):461–485, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020038.

Djaidja:2021:CBT

[405] Noui Djaidja and Mostefa Nadir. Comparison between Taylor and perturbed method for Volterra integral equation of the first kind. Numerical Algebra, Control and Optimization, 11(4):487–493, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020039.

Hoai:2021:AAS

[406] Nguyen Thi Hoai. Asymptotic approximation to a solution of a singularly perturbed linear-quadratic optimal control problem with second-order linear ordinary differential equation of state variable. Numerical Algebra, Control and Optimization, 11(4):495–512, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020040.

Hazzam:2021:PDI

[407] Nadia Hazzam and Zakia Kebbiche. A primal–dual interior point method for P_{∗} (κ)-HLCP based on a class of parametric kernel functions. Numerical Algebra, Control and Optimization, 11(4):513–531, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020053.

Mehrjerdi:2021:NMS

[408] Yahia Zare Mehrjerdi. A new methodology for solving bi-criterion fractional stochastic programming. Numerical Algebra, Control and Optimization, 11(4):533–554, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020054.

Karite:2021:GRC

[409] Touria Karite and Ali Boutoulout.

Global and regional constrained controllability for distributed parabolic linear systems: RHUM approach. Numerical Algebra, Control and Optimization, 11(4):555–566, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020055.

Khudher:2021:IBP

[410] Israa Mohammed Khudher, Yahya Ismail Ibrahim, and Suhaib Abduljabbar Altamir. Individual biometrics pattern based artificial image analysis techniques. Numerical Algebra, Control and Optimization, 11(4):567–578, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2020056.

Devarapalli:2021:NHA

[411] Ramesh Devarapalli and Biplab Bhattacharyya. A novel hybrid AGWO-PSO algorithm in mitigation of power network oscillations with STATCOM. Numerical Algebra, Control and Optimization, 11(4):579–611, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2020057.

Price:2021:MNM

[412] C. J. Price. A modified Nelder–Mead barrier method for constrained optimization. Numerical Algebra, Control and Optimization, 11(4):613–631, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL

                                        http://aimsciences.org/article/
                                        doi/10.3934/naco.2020058.

Qasim:2021:SND

[413] Omar Saber Qasim, Ahmed Entesar, and Waleed Al-Hayani. Solving nonlinear differential equations using hybrid method between Lyapunov’s artificial small parameter and continuous particle swarm optimization. Numerical Algebra, Control and Optimization, 11(4):633–644, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021001.

Aliane:2021:DMS

[414] Mohamed Aliane, Mohand Bentobache, Nacima Moussouni, and Philippe Marthon. Direct method to solve linearquadratic optimal control problems. Numerical Algebra, Control and Optimization, 11(4):645–663, ???? 2021. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021002.

Sun:2021:AGM

[415] Yanmei Sun and Yakui Huang. An alternate gradient method for optimization problems with orthogonality constraints. Numerical Algebra, Control and Optimization, 11(4):665–676, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021003.

Miao:2021:PIN

[416] Hong-Yi Miao and Li Wang. Preconditioned inexact Newton-like method for large nonsymmetric eigenvalue problems. Numerical Algebra, Control and Optimization, 11(4):677–685, ???? 2021. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021012.

Fang:2022:P

[417] Shu-Cherng Fang, Ruey-Lin Sheu, and Tamaki Tanaka. Preface. Numerical Algebra, Control and Optimization, 12(1):˘2170–˘2171, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021058.

Liao:2022:ARD

[418] Yu-Hsien Liao. Axiomatic results and dynamic processes for two weighted indexes under fuzzy transferable-utility behavior. Numerical Algebra, Control and Optimization, 12(1):1–14, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL

http://aimsciences.org/article/
doi/10.3934/naco.2021047.

Sun:2022:VFM

[419] Hsin-Min Sun and Yu-Juan Sun. Variable fixing method by weighted average for the continuous quadratic knapsack problem. Numerical Algebra, Control and Optimization, 12 (1):15–29, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021048.

Wang:2022:CDR

[420] I-Lin Wang and Chen-Tai Hou. A crowdsourced dynamic repositioning strategy for public bike sharing systems. Numerical Algebra, Control and Optimization, 12(1):31–46, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021049.

Miao:2022:LMM

[421] Xin-He Miao, Kai Yao, Ching-Yu Yang, and Jein-Shan Chen. Levenberg– Marquardt method for absolute value equation associated with second-order cone. Numerical Algebra, Control and Optimization, 12(1):47–61, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021050.

Kim:2022:WCT

[422] Do Sang Kim, Nguyen Ngoc Hai, and Bui Van Dinh. Weak convergence theorems for symmetric generalized hybrid mappings and equilibrium problems. Numerical Algebra, Control and Optimization, 12(1):63–78, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021051.

Hu:2022:MER

[423] Cheng-Feng Hu, Hsiao-Fan Wang, and Tingyang Liu. Measuring efficiency of a recycling production system with imprecise data. Numerical Algebra, Control and Optimization, 12(1):79–91, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021052.

Kerdkaew:2022:GOC

[424] Jutamas Kerdkaew, Rabian Wangkeeree, and Rattanaporn Wangkeeree. Global optimality conditions and duality theorems for robust optimal solutions of optimization problems with data uncertainty, using underestimators. Numerical Algebra, Control and Optimization, 12(1):93–107, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021053.

Zhao:2022:OPS

[425] Lianxia Zhao, Hui Qiao, and Qi An. Optimal pre-sale policy for deteriorating items. Numerical Algebra, Control and Optimization, 12(1):109–120, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021054.

Huang:2022:ODC

[426] Tone-Yau Huang and Tamaki Tanaka.

Optimality and duality for complex multi-objective programming. Numerical Algebra, Control and Optimization, 12(1):121–134, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021055.

Jiang:2022:ASS

[427] Canghua Jiang, Dongming Zhang, Chi Yuan, and Kok Ley Teo. An active set solver for constrained H_{∞} optimal control problems with state and input constraints. Numerical Algebra, Control and Optimization, 12(1):135–157, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021056.

Lin:2022:DRO

[428] Fengming Lin, Xiaolei Fang, and Zheming Gao. Distributionally Robust Optimization: a review on theory and applications. Numerical Algebra, Control and Optimization, 12(1):159–212, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021057.

Jin:2022:SNS

[429] Zhuo Jin, Ming Qiu, Ky Q. Tran, and George Yin. A survey of numerical solutions for stochastic control problems: Some recent progress. Numerical Algebra, Control and Optimization, 12(2):213–253, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2022004.

Owolabi:2022:NIP

[430] Abd semii Oluwatosin-Enitan Owolabi, Timilehin Opeyemi Alakoya, Adeolu Taiwo, and Oluwatosin Temitope Mewomo. A new inertial-projection algorithm for approximating common solution of variational inequality and fixed point problems of multivalued mappings. Numerical Algebra, Control and Optimization, 12(2):255–278, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021004.

Azi:2022:OCD

[431] Mourad Azi and Mohand Ouamer Bibi.

Optimal control of a dynamical system with intermediate phase constraints and applications in cash management. Numerical Algebra, Control and Optimization, 12(2):279–291, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021005.

Chakravarty:2022:CIA

[432] Jahnabi Chakravarty, Ashiho Athikho, and Manideepa Saha. Convergence of interval AOR method for linear interval equations. Numerical Algebra, Control and Optimization, 12 (2):293–308, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021006.

Swarnakar:2022:DTR

[433] Jaydeep Swarnakar. Discrete-time realization of fractional-order proportional integral controller for a class of fractional-order system. Numerical Algebra, Control and Optimization, 12(2):309–320, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021007.

Abed:2022:APS

[434] Abdulrazzaq T. Abed and Azzam S. Y. Aladool. Applying particle swarm optimization based on Padé approximant to solve ordinary differential equation. Numerical Algebra, Control and Optimization, 12(2):321–337, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021008.

Ge:2022:COS

[435] Zhaoqiang Ge. Controllability and observability of stochastic implicit systems and stochastic GE-evolution operator. Numerical Algebra, Control and Optimization, 12(2):339–351, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021009.

Mahmudov:2022:SOD

[436] Elimhan N. Mahmudov. Second order discrete time-varying and timeinvariant linear continuous systems and Kalman type conditions. Numerical Algebra, Control and Optimization, 12(2):353–371, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

                                        //aimsciences.org/article/doi/10.
                                        3934/naco.2021010.

Ogwo:2022:MEA

[437] Grace Nnennaya Ogwo, Chinedu Izuchukwu, and Oluwatosin Temitope Mewomo. A modified extragradient algorithm for a certain class of split pseudo-monotone variational inequality problem. Numerical Algebra, Control and Optimization, 12 (2):373–393, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021011.

Malmir:2022:CFD

[438] Iman Malmir. Caputo fractional derivative operational matrices of Legendre and Chebyshev wavelets in fractional delay optimal control. Numerical Algebra, Control and Optimization, 12(2):395–426, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021013.

Abdulaleem:2022:IDM

[439] Najeeb Abdulaleem. V -E-invexity in Edifferentiable multiobjective programming. Numerical Algebra, Control and Optimization, 12(2):427–443, ???? 2022. CODEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021014.

Faybusovich:2022:LSP

[440] Leonid Faybusovich and Cunlu Zhou.

Long-step path-following algorithm for quantum information theory: Some numerical aspects and applications. Numerical Algebra, Control and Optimization, 12(2):445–467, ???? 2022. CO-DEN ???? ISSN 2155-3289 (print), 2155-3297 (electronic). URL http:

//aimsciences.org/article/doi/10.
3934/naco.2021017.